### v1

# Noisy Timing Behavior is a Feature of Central Compact Object Pulsars

Abstract
We present a timing study of the three known central compact object (CCO) pulsars, isolated cooling neutron stars in supernova remnants, using Chandra, XMM-Newton and NICER observations spanning two decades. Relative to canonical young pulsars, CCOs are spinning down at a very slow rate $|\dot{f}|<10^{-15}$ s-2, implying a surface dipole magnetic field strength $B_{s}<10^{11}$ G that is too weak to account for their X-ray emitting hot spots. Two CCO pulsars with sufficiently long monitoring, 1E 1207.4-5209 and PSR J0821-4300, are seen to deviate from steady spin-down; their timing residuals can be modeled by one or more glitches in $f$ and $\dot{f}$, or alternatively by extreme timing noise. For the third CCO pulsar, PSR J1852++0040, the sparse temporal coverage was insufficient to detect such effects. Glitch activity and timing noise in large samples of rotation-powered pulsars correlate best with $\dot{f}$, while the timing irregularities of the first two CCOs are extreme compared to pulsars of the same $\dot{f}$. Nevertheless, timing activity in CCOs may arise from properties that they share with other young but more energetic pulsars: high internal temperature, strong buried magnetic field and superfluid behavior. Alternatively, continuing low-level accretion of supernova debris is not ruled out as a source of timing noise in CCOs.
Unified Astronomy Thesaurus concepts: Pulsars (1306); Neutron stars (1108); Compact objects (288); Pulsar timing method (1305)
1 Introduction
Central compact objects (CCOs) comprise a dozen or so cooling neutron stars (NSs) in supernova remnants (SNRs), defined and distinguished from canonical, young rotation-powered pulsars by their steady surface thermal X-ray emission, lack of surrounding pulsar wind nebula, and non-detection at any other wavelength (Pavlov et al. 2002; see De Luca 2017 for a review). Considering that their faintness and spectral softness are obstacles to widespread discovery in the Galactic plane, CCOs may be as common as more conspicuous NS classes, e.g., magnetars, and may represent a major channel of NS birth.
Pulsations have been detected in only three of the CCOs: 1E 1207.4-5209 in SNR PKS 1209-51/52, PSR J0821-4300 in SNR Puppis A, and PSR J1852++0040 in SNR Kes 79. They have typical periods ($0.424,0.112,0.104$ s) but small spin-down rates $\dot{f}=(-1.2,-6.8,-7.9)\times 10^{-16}$ s-2, implying surface dipole magnetic field strengths of $(9.8,2.9,3.1)\times 10^{10}$ G, respectively, which are exceptionally weak for such young pulsars (Gotthelf et al., 2013; Halpern & Gotthelf, 2010). Furthermore, the surface magnetic field strength of 1E 1207.4-5209 inferred from spin-down, $B_{s}=9.8\times 10^{10}$ G, is close to $B\approx 8\times 10^{10}$ G, the value measured from its unique series of spectroscopic absorption features (Bignami et al., 2003), interpreted as the electron cyclotron fundamental at 0.7 keV and its harmonics. The similar values of field strength obtained by the two independent methods is convincing evidence of the weak surface $B$ field for at least this one CCO.
The remainder of the CCOs, while having similar continuum spectral properties as the CCO pulsars, have eluded searches for pulsations (Alford & Halpern, 2023). It is postulated that they may have even weaker magnetic fields, more uniform surface temperature distributions, or an unfavorable viewing geometry. In fact, it is difficult to create the hot spots needed to produce the X-ray pulsations observed in CCOs assuming only their weak surface dipoles. The only mechanism thought to be capable of creating a nonuniform surface temperature is anisotropic heat conduction in a strong magnetic field. As recounted in Gotthelf & Halpern (2018), models that invoke strong non-dipolar fields in the crust have long been studied. More recently, e.g., Igoshev et al. (2021) calculated the magnetothermal evolution of strong crustal fields generated by a stochastic dynamo, finding a tangled geometry with negligible dipole component. This creates factors of  2 difference in surface temperature on small scales, sufficient to produce the observed pulsations in CCOs.
Postulating the existence of strong non-dipole crustal fields was also the key to an explanation for the conspicuous absence of CCO descendants, the large population of weak $B$-field pulsars that should remain in the same location of the $P-\dot{P}$ diagram after their SNRs have faded. If CCOs are actually born with a canonical NS magnetic dipole field that was promptly buried by fallback of a small amount $(\sim 10^{-4}\,M_{\odot})$ of supernova debris, the buried field will diffuse back to the surface on a timescale of $\sim 10^{5}$ years (Muslimov & Page, 1995; Bernal et al., 2010; Ho, 2011; Viganò & Pons, 2012). A CCO will move vertically up in the $P-\dot{P}$ diagram as its dipole field grows back (Bogdanov et al., 2014; Ho, 2015; Luo et al., 2015), merging with the population of ordinary radio pulsars.
The discovery of a glitch and/or strong timing noise from 1E 1207.4-5209 (Gotthelf & Halpern, 2018, 2020) was of timely relevance to these theories of CCOs. Glitches are thought to result from either “starquakes”, stress relief of the NS crust, or from unpinning of superfluid vortices in the inner part of the NS crust (e.g., Link et al., 1998). A review of glitch observations and theory is given by Antonopoulou et al. (2022), while a comprehensive compilation of glitch statistics can be found in Espinoza et al. (2011) and Fuentes et al. (2017). Timing noise has been attributed to variability in the interaction of the crustal superfluid with the Coulomb lattice of the solid crust (Jones, 1990), turbulence of the superfluid (Melatos & Link, 2014), or fluctuations in the structure of the magnetosphere, e.g., state switching (Lyne et al., 2010). Analysis methods for timing noise are reviewed in Parthasarathy et al. (2019) and Namkham et al. (2019).
The timing irregularities in 1E 1207.4-5209 were extreme relative to the general pulsar population, in which glitch activity is correlated primarily with frequency derivative $\dot{f}$. Naively, glitches are unexpected in CCOs because of their small $\dot{f}$. However, the idea of stronger internal $B$ fields, needed to explain CCO surface hot spots and population statistics, may have an additional application in producing the timing irregularities of CCOs. Ho (2015) proposed that glitch activity could be triggered by the motion of strong magnetic fields through the NS crust, interacting with the neutron superfluid there, and that glitches would be a way of identifying CCO descendants.
To add complexity, it has not been ruled out that an entirely different process, low-level accretion in the propeller regime, could cause the observed timing noise, specifically the observed changes in $\dot{f}$ (Halpern et al., 2007; Gotthelf & Halpern, 2018). This, because the required accretion rate of $\leq 10^{-11}$ g s-1 makes a negligible contribution to the observed luminosity, while the mass supply could be a residual fallback disk of $\sim 10^{-6}\,M_{\odot}$, much less mass than may have fallen back on the NS initially. The spin parameters of CCOs fall in regime where dipole braking and accretion torque noise could be comparable.
The subject of this paper is the long-term spin-down evolution of all three CCO pulsars spanning the two decades since their discovery. In Section 2 we present new Chandra, XMM-Newton, and NICER timing observations which we use to update timing solutions, explore alternative timing models, and identify timing irregularites in PSR J0821-4300 for the first time. Section 3 discusses the implications of the candidate glitches and timing noise, and compares the CCO results with those of the general pulsar population. Section 4 summarizes our conclusions.
2 Timing Analysis
We obtained new timing observations for the three CCO pulsars as part of our continued monitoring programs using the Chandra and XMM-Newton observatories. We were able to achieve better temporal coverage by requesting observations from the two missions, allowing for their unique scheduling constraints. With the onset of the NICER mission, we continued our coverage of 1E 1207.4-5209 using NICER exclusively. The previously published data sets also included here are fully described in our earlier work (1E 1207.4-5209 — Gotthelf & Halpern, 2007; Halpern & Gotthelf, 2011; Gotthelf et al., 2013; Halpern & Gotthelf, 2015; Gotthelf & Halpern, 2018, 2020), (PSR J0821-4300 — Gotthelf & Halpern, 2009; Gotthelf et al., 2013), and (PSR J1852++0040 — Gotthelf et al., 2005; Halpern et al., 2007; Halpern & Gotthelf, 2010).
All data sets were reprocessed and reduced using the latest software for each mission, following the methods described in our earlier papers. Photon arrival times were converted to barycentric dynamical time (TDB) using the DE405 solar system ephemeris and the Chandra coordinates given in the tables below for each object. For PSR J0821-4300, we account for its high proper motion using the ephemeris of Mayer et al. (2020). Photons were extracted using a circular aperture optimized for each target and instrument, based on the local SNR contamination in the aperture.
For each observation, we generated a phase zero time-of-arrival (ToA) measurement to use in determining the long-term spin evolution of the NS, as follows. Starting from the filtered event files, we measured the pulse period at each epoch using a periodogram search for the maximum power around the expected frequency, as informed by our earlier work. For 1E 1207.4-5209 and PSR J0821-4300, their pulse profiles are well-characterized by a sine function, and we computed phase zero from the Fourier coefficients of the unbinned photon arrival times. We define phase zero as the peak of the sine function (see Gotthelf & Halpern, 2020, for details). For PSR J1852++0040, which has a much broader pulse profile, a high-significance pulse profile template was generated by adding all the folded profiles, aligned by cross-correlating with an iterated template. We define phase zero for this CCO as the middle of the valley in the pulse profile.
Utilizing the TEMPO software (Hobbs et al., 2006), we initiated a phase-connected timing solution to a subset of ToAs for each pulsar, using a model for the rotation phase of the pulsar that includes one or two of its frequency derivatives:
$\phi(t)=\phi_{o}+f(t-t_{o})+{1\over 2}\dot{f}(t-t_{o})^{2}+{1\over 6}\ddot{f}(t-t_{o})^{3}.$
Generally, we obtained optimally spaced sets of observations that allow us to start and to maintain a phase-connected timing solution that accounts for every turn of the pulsar, with increased accuracy, by bootstrapping to longer intervals. This is usually possible unless the interval to the next ToA is too large to predict its cycle count by less that $\sim 10\%$ in phase. As described below, due to data gaps and timing irregularities, it was necessary at some point to restart a new solution for each pulsar to continue to follow and/or interpret its spin-down behavior. However, if the timing residuals begin to deviate significantly and systematically from the trial solution, it signifies that the ephemeris no longer predicts the pulsar’s spin-down, and more complex models need to be considered.
2.1 1E 1207.4-5209
In this work we use all available data sets for 1E 1207.4-5209 from XMM-Newton, Chandra, and NICER, as logged in Table 1. These span 23 yr and include 3.5 yr of monitoring since our previously published results (Gotthelf & Halpern, 2020). The new observations used here were acquired exclusively using NICER. In the following, we extend the post-2015 glitch ephemeris as well as update the alternative models suggested by the earlier work. We do not repeat the previously reported fit for a possible pre-2015 glitch from 1E 1207.4-5209, as that result is unchanged. In the following analysis, we limit the photon energy range to 0.5–1.6 keV, optimal for the pulsar’s observed soft X-ray spectrum.
[…]
Table 1: Log of X-ray Timing Observations of 1E 1207.4-5209
Mission Instrument ObsIDaaThe NICER data sets are denoted by the ObsID and date of the
first of the concatenated set of short adjacent observations.
Date ExpbbExposure times for the XMM-Newton EPIC-pn do not reflect the
29% deadtime in SmallWindow (sw) mode.
/Mode (UT) (ks)
Chandra ACIS-S3/$C$ 751 2000 Jan 6 32.4
XMM EPIC-pn/sw 0113050501 2001 Dec 23 27.0
Chandra ACIS-S/$C$ 2799 2002 Jan 5 30.3
XMM EPIC-pn/sw 0155960301 2002 Aug 4 128.0
XMM EPIC-pn/sw 0155960501 2002 Aug 6 129.0
Chandra ACIS-S/$C$ 3915 2003 Jun 10 155.1
Chandra ACIS-S/$C$ 4398 2003 Jun 18 114.7
XMM EPIC-pn/sw 0304531501 2005 Jun 22 15.1
XMM EPIC-pn/sw 0304531601 2005 Jul 5 18.2
XMM EPIC-pn/sw 0304531701 2005 Jul 10 20.5
XMM EPIC-pn/sw 0304531801 2005 Jul 11 63.4
XMM EPIC-pn/sw 0304531901 2005 Jul 12 14.5
XMM EPIC-pn/sw 0304532001 2005 Jul 17 16.5
XMM EPIC-pn/sw 0304532101 2005 Jul 31 17.7
XMM EPIC-pn/sw 0552810301 2008 Jul 2 31.4
XMM EPIC-pn/sw 0552810401 2008 Dec 22 30.4
Chandra ACIS-S3/$C$ 14199 2011 Nov 25 31.0
Chandra ACIS-S3/$C$ 14202 2012 Apr 10 33.0
XMM EPIC-pn/sw 0679590101 2012 Jun 22 26.5
XMM EPIC-pn/sw 0679590201 2012 Jun 24 22.3
XMM EPIC-pn/sw 0679590301 2012 Jun 28 24.9
XMM EPIC-pn/sw 0679590401 2012 Jul 2 24.5
XMM EPIC-pn/sw 0679590501 2012 Jul 18 27.3
XMM EPIC-pn/sw 0679590601 2012 Aug 11 27.3
Chandra ACIS-S3/$C$ 14200 2012 Dec 1 31.1
Chandra ACIS-S3/$C$ 14203 2013 May 19 33.0
Chandra ACIS-S3/$C$ 14201 2013 Dec 4 33.0
Chandra ACIS-S3/$C$ 14204 2014 Jun 20 33.0
XMM EPIC-pn/sw 0780000201 2016 Jul 28 32.5
XMM EPIC-pn/sw 0800960201 2017 Jun 22 33.3
XMM EPIC-pn/sw 0800960301 2017 Jun 23 20.7
XMM EPIC-pn/sw 0800960401 2017 Jun 24 22.6
XMM EPIC-pn/sw 0800960501 2017 Jul 3 23.5
NICER XTI 1020270102 2017 Jul 24 6.1
NICER XTI 1020270106 2017 Jul 28 14.3
NICER XTI 1020270110 2017 Aug 1 12.1
XMM EPIC-pn/sw 0800960601 2017 Aug 10 19.8
Chandra ACIS-S3/$C$ 19612 2017 Oct 10 32.9
NICER XTI 1020270130 2017 Nov 15 20.6
XMM EPIC-pn/sw 0800960701 2017 Dec 24 19.8
XMM EPIC-pn/sw 0821940201 2018 Jun 22 33.2
Chandra ACIS-S3/$C$ 19613 2018 Aug 27 66.3
NICER XTI 1020270153-58 2018 Nov 30 7.6
XMM EPIC-pn/sw 0821940301 2018 Dec 28 26.6
NICER XTI 2506010101-02 2019 Apr 4 22.3
XMM EPIC-pn/sw 0842280301 2019 Jul 9 30.8
NICER XTI 2506010201-02 2019 Jul 19 21.4
NICER XTI 2506010205-13 2019 Jul 26 6.7
New Observations – This Work
NICER XTI 2506010214-24 2019 Aug 29 10.6
NICER XTI 2506010301-03 2019 Oct 31 15.6
NICER XTI 2506010304-409 2019 Nov 19 9.8
NICER XTI 2506010415-20 2020 Feb 16 19.0
NICER XTI 3550010101-03 2020 Mar 30 17.8
NICER XTI 3550010201-03 2020 Jun 12 18.6
NICER XTI 3550010401-03 2020 Dec 23 17.7
NICER XTI 3550010501-02 2021 Feb 27 17.3
NICER XTI 4614010101-03 2021 Apr 4 16.3
NICER XTI 4614010201-04 2021 May 28 14.5
NICER XTI 4614010301-04 2021 Jun 29 12.9
NICER XTI 4614010501-02 2021 Nov 19 8.2
NICER XTI 4614010601-02 2022 Jan 1 8.5
NICER XTI 4614010701-02 2022 Feb 4 17.4
NICER XTI 4614010801-03 2022 Apr 19 17.7
NICER XTI 4614010901-05 2022 Jun 16 21.1
NICER XTI 4614011001-02 2022 Sep 1 8.0
NICER XTI 4614011201-03 2023 Feb 11 15.8
Table 2: Glitch Ephemerides for 1E 1207.4-5209footnotetext: Notes. Derived parameters ($B_{s}$, $\dot{E}$, $\tau_{c}$) are based on the pre-glitch timing solution.
[…]
R.A. (J2000) $12^{\rm h}10^{\rm m}00^{\rm s}\!.91$
Decl. (J2000) $52^{\circ}26^{\prime}28^{\prime\prime}\!.4$
Surface dipole magnetic field, $B_{s}$ $9.8\times 10^{10}$ G
Spin-down luminosity, $\dot{E}$ $1.1\times 10^{31}$ erg s-1
Characteristic age, $\tau_{c}\equiv P/2\dot{P}$ 303 Myr
Pre-glitch Timing Solution (2002–2014)
Epoch of ephemeris (MJD TDB) 54547.00000198
Span of ephemeris (MJD) 52266–56829
Frequency, $f$ 2.357763492491(28) s-1
Frequency derivative, $\dot{f}$ $1.2317(66)\times 10^{-16}$ s-2
Period, $P$ 0.4241307506816(50) s
Period derivative, $\dot{P}$ $2.216(12)\times 10^{-17}$
$\chi^{2}_{\nu}[{\rm DoF}]$ 1.80[25]
Post-Glitch Timing Solution (2016–2023)
Epoch of ephemeris (MJD TDB) 58144.00000219
Span of ephemeris (MJD) 57597–59986
Frequency, $f$ 2.35776345898(14) s-1
Frequency derivative, $\dot{f}$ $1.120(27)\times 10^{-16}$ s-2
Period, $P$ 0.424130756710(26) s
Period derivative, $\dot{P}$ $2.015(49)\times 10^{-17}$
$\chi^{2}_{\nu}[{\rm DoF}]$ 1.26[35]
Glitch epoch (MJD) 57041bbEpoch of the glitch estimated by matching the frequency of the two timing solutions; this assumes a constant post-glitch $\dot{f}$.
$\Delta f$ $3.87(39)\times 10^{-9}$ s-1
$\Delta f/f_{\rm pred}$ $1.64(16)\times 10^{-9}$
$\Delta\dot{f}$ $1.12(28)\times 10^{-17}$ s-2
$\Delta\dot{f}/\dot{f}$ 0.091(23)-0.091(23)
Table 3: Alternative Timing Solutions for 1E 1207.4-5209
[…]
Quadratic Timing Solution (2000–2023)
Epoch of ephemeris (MJD TDB) 55478.00000467
Span of ephemeris (MJD) 51549–59986
Frequency, $f$ 2.357763483778(12) s-1
Frequency derivative, $\dot{f}$ $1.1128(15)\times 10^{-16}$ s-2
Period, $P$ 0.4241307522490(21) s
Period derivative, $\dot{P}$ $2.0018(28)\times 10^{-17}$
$\chi^{2}_{\nu}[{\rm DoF}]$ 8.55[64]
Cubic Timing Solution (2000–2023)
Epoch of ephemeris (MJD TDB) 55478.00000453
Span of ephemeris (MJD) 51549–59986
Frequency, $f$ 2.357763483219(32) s-1
Frequency derivative, $\dot{f}$ $1.1260(17)\times 10^{-16}$ s-2
Frequency second derivative, $\ddot{f}$ $4.67(25)\times 10^{-26}$ s-3
Period, $P$ 0.4241307523495(58) s
Period derivative, $\dot{P}$ $2.0255(30)\times 10^{-17}$
Period second derivative, $\ddot{P}$ $8.40(45)\times 10^{-27}$ s-1
$\chi^{2}_{\nu}[{\rm DoF}]$ 3.17[63]
Cubic Timing Solution (2002–2023)
Epoch of ephemeris (MJD TDB) 55478.00000447
Span of ephemeris (MJD) 52266–59986
Frequency, $f$ 2.357763483142(35) s-1
Frequency derivative, $\dot{f}$ $1.1318(20)\times 10^{-16}$ s-2
Frequency second derivative, $\ddot{f}$ $5.51(29)\times 10^{-26}$ s-3
Period, $P$ 0.42413075236(63) s
Period derivative, $\dot{P}$ $2.0359(36)\times 10^{-17}$
Period second derivative, $\ddot{P}$ $0.92(53)\times 10^{-26}$ s-1
$\chi^{2}_{\nu}[{\rm DoF}]$ 2.72[62]
Figure 1: Glitch model for the spin-down of 1E 1207.4-5209. Top: pulse-phase residuals from the pre-glitch timing solution presented in Table 2. The glitch epoch of 2015 January 19 is estimated by matching the frequency of the pre- and post-glitch solutions. The year 2000 Chandra data point is not included in the fit (see Section 2.1 for details). Bottom: combined residuals from fits to independent timing models for the pre- and post-glitch intervals, respectively. The overall $\chi^{2}_{\nu}$ statistic for the fit is 1.49 for 60 DoF, taking into account the fit parameters for each interval.
Figure 2: Pulse-phase residuals for 1E 1207.4-5209 using the alternative timing models presented in Table 3 that do not involve a glitch. These fits include the year 2000 Chandra data point (see Section 2.1 for details). Top: A quadratic model fit leaves a sinusoidal-like oscillation in the phase residuals. Bottom: the residuals from a cubic model fit (including the frequency second derivative).
Figure 1 graphs the ToA phase residuals from the pre-glitch timing solution given in Table 2, developed using data points from 2002–2014 (as in Gotthelf & Halpern, 2020). As noted in our earlier work, the very first ToA, the Chandra observation of 2000, does not appear consistent with the pre-glitch timing solution, and may indicate an even earlier glitch. We again excluded this data point from the fit, but continue to plot it in Figure 1 for reference.
After 2015, deviations from the expected phase of this solution were evident, and we were no longer able to extend the phase-connected solution forward in time. Instead, we postulated a glitch, and found that the subsequent ToAs could be characterized by a new, post-glitch phase-connected ephemeris. In this work, we have extended the post-glitch ephemeris, which remains consistent with the earlier result and shows no evidence of a newer glitch. The updated change of frequency between the post- and pre-glitch timing solutions is $\Delta f=(3.87\pm 0.39)\times 10^{-9}$ Hz, with a glitch magnitude of $\Delta f/f_{\rm pred}=(1.64\pm 0.16)\times 10^{-9}$. The post-glitch solution matches $f_{\rm pred}$, the frequency predicted by the pre-glitch solution on 2015 January 19 (MJD 57041), which we estimate as the epoch of the glitch (ignoring any short-timescale recovery behavior, which would not have been sampled). Of note, the glitch magnitude for 1E 1207.4-5209 is typical of rotation-powered pulsars with strong magnetic fields and rapid spin-down, including the Crab. For reference, $\Delta f/f=3\times 10^{-9}$ is where the lower-amplitude peak in the bimodal distribution of glitch magnitudes is centered (Espinoza et al. 2011).
The magnitude of the glitch is 23% smaller than the value quoted in Gotthelf & Halpern (2020). Nevertheless, this is the same post-glitch timing solution as the one published previously in the sense that the cycle counts calculated over the common time span are the same. The fit statistic of the combined pre- and post-glitch data sets is $\chi^{2}_{\nu}=1.49$ for 60 DoF. This is to be compared to $\chi^{2}_{\nu}=1.44$ for 42 DoF reported previously for the shorter span.
The uncertainty on the frequency derivative of the revised post-glitch timing solution has grown, perhaps an effect of timing noise. In addition, the post-glitch frequency derivative is larger (less negative) than the pre-glitch value by $\approx 9\%$, albeit at the $4\sigma$ level. This change is much larger than the $\Delta\dot{f}/\dot{f}\sim 10^{-3}$ typically seen across glitches, which also suggests that timing noise may predominate.
Weak glitches can be difficult to detect in the presence of timing noise, and timing noise can masquerade as glitches (Antonopoulou et al., 2022). Following Gotthelf & Halpern (2020), we therefore considered alternative timing models that do not require the use of a glitch to characterize the timing behavior of 1E 1207.4-5209. First, we show that a fit to the data using a quadratic model only, parameterized by the frequency and first derivative, and including the year 2000 observation, leaves a strong sinusoidal-like residual in the phases, with an amplitude of $\sim 0.1$ cycles (see Figure 2). As in Gotthelf & Halpern (2020), this results in an unacceptable fit statistic of $\chi^{2}_{\nu}=8.549$ for 64 DoF, substantially worse than for the glitch model. The sinusoidal residual with a period similar to the data span suggests that a polynomial term of higher degree should be added to the model. The lower panel of Figure 2 shows the result of fitting a cubic model that includes a frequency second derivative. This reduces the statistic to $\chi^{2}_{\nu}=3.17$ for 63 DoF, still significantly higher than for the glitch model. By excluding the 2000 data point, the fit statistic improves marginally to $\chi^{2}_{\nu}=2.72$ for 62 DoF, but still suggests larger than expected timing noise, as will be discussed in Section 3.
Parameters from the quadratic and cubic fits are given in Table 3. It is important that each of these models, as well as the glitch model, have the same cycle count between all adjacent ToAs. Thus they are all the same phase-connected solution that we believe is the true one, not an alias. Finally, we note that the binary orbit model suggested in Gotthelf & Halpern (2020) remains a viable interpretation, although less likely a priori.
2.2 PSR J0821-4300
Table 4 is the log of timing observations of PSR J0821-4300 used in this work. These include the Chandra and XMM-Newton data sets previously reported in Gotthelf et al. (2013) that span years 200920122009-2012, but exclude the two observations from 2001. In retrospect, it is not possible to span the large time gap from 2001 to 2009 due to the uncertainty in extrapolation of the pre-glitch ephemeris. In creating ToAs for PSR J0821-4300, we restrict the extracted photon energies to the 1.54.51.5-4.5 keV range because of the unique energy dependence of the phase of the pulse. Below 1.5 keV, the pulse is shifted $\approx 180^{\circ}$, which effectively cancels the pulsations if the full energy band is used (see Gotthelf & Halpern, 2009, for details). Photon arrival times were extracted using $30^{\prime\prime}$ and $1\farcs 2$ radius circular apertures for the XMM-Newton and Chandra data sets, respectively.
Following the results presented in Gotthelf et al. (2013), we were able to continue monitoring this pulsar and extend its ephemeris, but only up to 2014. Subsequently, the yearly monitoring showed increasing departures of the pulse arrival times from the refined ephemeris. Motivated by the similarity to the glitch-like behavior of 1E 1207.4-5209, we then initiated a renewed joint XMM-Newton and Chandra timing campaign on PSR J0821-4300 in 201920202019-2020 to generate an independent phase-connected timing solution for comparison.
In Figure 3, we postulate that a glitch in the pulse arrival times most likely occurred between the 2015 and 2016 observations. Indeed, the monotonic deviations from the pre-glitch timing solution given in Table 5 imply a change in frequency of magnitude $\Delta f/f_{\rm pred}=7.1(6)\times 10^{-10}$, a value similar to that found for 1E 1207.4-5209. Considering the pre- and post-glitch timing solutions, the statistic for the combined data set is $\chi^{2}_{\nu}=2.20$ for 32 DoF, taking into account the fit parameters for each interval.
The sparse coverage around the time of the glitch, between 2014 and 2017, prevents a definitive description of the glitch behavior. However, the formal fits show curvature in the trend of the post-glitch residuals relative to the pre-glitch solution, corresponding to a $\approx 7\%$ change in frequency derivative, becoming less negative, similar to the case of 1E 1207.4-5209 but at a higher $9\sigma$ level of significance (see Table 5).
Curiously, a model involving an anti-glitch (decrease in frequency) fits nearly as a well as the glitch model, with the event occurring at an earlier time, around 2011 September (Figure 4). This model parameters are given in Table 6, where it can be seen that magnitude of the anti-glitch is similar to that in the glitch model, while the statistic for the combined data set is slightly worse at $\chi^{2}_{\nu}=2.77$. Since the pre-glitch segment in this model is very short, the initial frequency derivative was only measured to a precision of $\approx 8\%$, but it is consistent with the post-glitch value. Note that this model has the same set of cycle counts as the glitch model, despite being a different analytic and qualitative description of the same timing data.
We also consider alternative timing models without the use of a glitch to characterize the timing of PSR J0821-4300. A basic quadratic model leaves phase residuals with $\sim 0.1$ cycle amplitude, similar in magnitude to those of 1E 1207.4-5209, but not clearly sinusoidal in nature (Figure 5). The fit statistic is $\chi^{2}_{\nu}=6.79$ for 35 DoF, again substantially worse than for the glitch model and similar to that found for 1E 1207.4-5209. The additional of a second frequency derivative term for a cubic model has little effect on the shape of the residual curve, and only reduces $\chi^{2}_{\nu}$ to 5.80 for 34 DoF. Parameters from the quadratic and cubic fits are given in Table 7. It can be seen how the residual curve in Figure 5 could also be interpreted as an early anti-glitch or a later glitch. Again, none of these are distinct timing solutions (aliases); they all have the same set of cycle counts.
Table 4: Log of X-ray Timing Observations of PSR J0821-4300
Mission Instrument ObsID Date ExpaaExposure times for the XMM-Newton EPIC-pn do not reflect the 29% deadtime in SmallWindow (sw) mode.
/Mode (UT) (ks)
XMM EPIC-pn/sw 0606280101 2009 Dec 17, 18 85
XMM EPIC-pn/sw 0606280201 2010 Apr 5 42
XMM EPIC-pn/sw 0650220201 2010 May 2 28
Chandra ACIS-S3/$C$ 12108 2010 Aug 16 34
XMM EPIC-pn/sw 0650220901 2010 Oct 15 24
XMM EPIC-pn/sw 0650221001 2010 Oct 15 24
XMM EPIC-pn/sw 0650221101 2010 Oct 19 27
XMM EPIC-pn/sw 0650221201 2010 Oct 25 25
XMM EPIC-pn/sw 0650221301 2010 Nov 12 24
XMM EPIC-pn/sw 0650221401 2010 Dec 20 27
Chandra ACIS-S3/$C$ 12109 2011 Feb 4 33
XMM EPIC-pn/sw 0650221501 2011 Apr 12 30
XMM EPIC-pn/sw 0657600101 2011 May 18 37
Chandra ACIS-S3/$C$ 12541 2011 Aug 11 33
XMM EPIC-pn/sw 0657600201 2011 Nov 8 37
Chandra ACIS-S3/$C$ 12542, 14395 2012 Feb 18, 19 33
XMM EPIC-pn/sw 0657600301 2012 Apr 10 35
New Observations – This Work
Chandra ACIS-S3/$C$ 14798 2012 Dec 11 31
XMM EPIC-pn/sw 0722640301 2013 Oct 29 45
XMM EPIC-pn/sw 0722640401 2013 Oct 31 42
Chandra ACIS-S3/$C$ 16254 2014 Jul 21 33
XMM EPIC-pn/sw 0742040201 2014 Oct 18 46
Chandra ACIS-S3/$C$ 16255 2015 Sep 7 33
Chandra ACIS-S3/$C$ 16256 2016 Sep 18 34
XMM EPIC-pn/sw 0781870101 2016 Nov 8 91
Chandra ACIS-S3/$C$ 19609 2017 Aug 14 12
Chandra ACIS-S3/$C$ 19610,21110 2018 Jun 26, 28 34
Chandra ACIS-S3/$C$ 19611 2019 Aug 2 33
XMM EPIC-pn/sw 0853220201 2019 Oct 9 23
XMM EPIC-pn/sw 0853220301 2019 Oct 11 26
XMM EPIC-pn/sw 0853220401 2019 Oct 16 34
XMM EPIC-pn/sw 0853220501 2019 Nov 10 26
XMM EPIC-pn/sw 0853220601 2019 Dec 26 23
Chandra ACIS-S3/$C$ 22674 2020 Aug 17 45
XMM EPIC-pn/sw 0882950101 2021 May 3 37
Chandra ACIS-S3/$C$ 22675, 25984 2021 Aug 25, 27 48
XMM EPIC-pn/sw 0882950201 2021 Oct 22 35
Chandra ACIS-S3/$C$ 22676, 26477 2022 Jul 27, 28 52
Table 5: Glitch Ephemerides for PSR J0821-4300footnotetext: Notes. Position and proper motion from Mayer et al. (2020); Derived parameters ($B_{s}$, $\dot{E}$, $\tau_{c}$) are based on the pre-glitch timing solution.
[…]
Epoch of position and $\mu$ (MJD) 53964.0
R.A. (J2000) $08^{\rm h}21^{\rm m}57.\!^{\rm s}3653(31)$
Decl. (J2000) $43^{\circ}00^{\prime}17.\!^{\prime\prime}074(43)$
R.A. proper motion, $\mu_{\alpha}\,{\rm cos}\,\delta$ $54.1\pm 8.3$ mas yr-1
Decl. proper motion, $\mu_{\delta}$ $28.1\pm 10.5$ mas yr-1
Surface dipole magnetic field, $B_{s}$ $2.9\times 10^{10}$ G
Spin-down luminosity, $\dot{E}$ $1.9\times 10^{32}$ erg s-1
Characteristic age, $\tau_{c}$ 254 Myr
Pre-glitch Timing Solution (2010–2014)
Epoch of ephemeris (MJD TDB) 56065.00000030
Span of ephemeris (MJD) 55182–56948
Frequency, $f$ 8.865291025552(85) s-1
Frequency derivative, $\dot{f}$ $6.801(44)\times 10^{-16}$ s-2
Period, $P$ 0.1127994554401(11) s
Period derivative, $\dot{P}$ $8.654(56)\times 10^{-18}$
$\chi^{2}_{\nu}[{\rm DoF}]$ 2.70[19]
Post-Glitch Timing Solution (2015–2022)
Epoch of ephemeris (MJD TDB) 58530.00000022
Span of ephemeris (MJD) 57272–59787
Frequency, $f$ 8.865290891814(75) s-1
Frequency derivative, $\dot{f}$ $6.321(27)\times 10^{-16}$ s-2
Period, $P$ 0.11279945714171(95) s
Period derivative, $\dot{P}$ $8.042(34)\times 10^{-18}$
$\chi^{2}_{\nu}[{\rm DoF}]$ 1.47[13]
Glitch epoch (MJD) 57362bbEpoch of the glitch estimated by matching the frequency of the two timing solutions; this assumes a constant post-glitch $\dot{f}$.
$\Delta f$ $6.26(55)\times 10^{-9}$ s-1
$\Delta f/f_{\rm pred}$ $7.06(62)\times 10^{-10}$
$\Delta\dot{f}$ $4.80(52)\times 10^{-17}$ s-2
$\Delta\dot{f}/\dot{f}$ 0.0706(76)-0.0706(76)
Table 6: Anti-glitch Ephemerides for PSR J0821-4300bbfootnotetext: Epoch of the glitch estimated by matching the frequency of the two timing solutions; this assumes a constant post-glitch $\dot{f}$.
[…]
Pre-glitch Timing Solution (2010–2011)
Epoch of ephemeris (MJD TDB) 55484.00000014
Span of ephemeris (MJD) 55182–55784
Frequency, $f$ 8.86529106058(42) s-1
Frequency derivative, $\dot{f}$ $6.14(50)\times 10^{-16}$ s-2
Period, $P$ 0.1127994549943(53) s
Period derivative, $\dot{P}$ $7.81(64)\times 10^{-18}$
$\chi^{2}_{\nu}[{\rm DoF}]$ 2.535[11]
Post-Glitch Timing Solution (2012–2022)
Epoch of ephemeris (MJD TDB) 57830.00000060
Span of ephemeris (MJD) 55873–59787
Frequency, $f$ 8.865290929387(33) s-1
Frequency derivative, $\dot{f}$ $6.2179(81)\times 10^{-16}$ s-2
Period, $P$ 0.11279945666364(42) s
Period derivative, $\dot{P}$ $7.912(10)\times 10^{-18}$
$\chi^{2}_{\nu}[{\rm DoF}]$ 2.695[21]
Glitch epoch (MJD) 55828b
$\Delta f$ $5.4(6)\times 10^{-9}$ s-1
$\Delta f/f_{\rm pred}$ $6.1(6)\times 10^{-10}$
Table 7: Alternative Timing Solutions for PSR J0821-4300
[…]
Quadratic Timing Solution (2010–2022)
Epoch of ephemeris (MJD TDB) 57485.00000096
Span of ephemeris (MJD) 55182–59787
Frequency, $f$ 8.865290948152(24) s-1
Frequency derivative, $\dot{f}$ $6.2723(51)\times 10^{-16}$ s-2
Period, $P$ 0.11279945642488(30) s
Period derivative, $\dot{P}$ $7.9807(65)\times 10^{-18}$
$\chi^{2}_{\nu}[{\rm DoF}]$ 6.79[35]
Cubic Timing Solution (2010–2022)
Epoch of ephemeris (MJD TDB) 57485.00000097
Span of ephemeris (MJD) 55182–59787
Frequency, $f$ 8.865290947727(71) s-1
Frequency derivative, $\dot{f}$ $6.2665(52)\times 10^{-16}$ s-2
Frequency second derivative, $\ddot{f}$ $1.00(16)\times 10^{-25}$ s-3
Period, $P$ 0.11279945643030(91) s
Period derivative, $\dot{P}$ $7.9733(66)\times 10^{-18}$
Period second derivative, $\ddot{P}$ $1.27(20)\times 10^{-27}$ s-1
$\chi^{2}_{\nu}[{\rm DoF}]$ 5.81[34]
Figure 3: Glitch model for the spin-down of PSR J0821-4300. Top: pulse-phase residuals from the pre-glitch timing solution presented in Table 5. The glitch epoch of 2015 December 6 is estimated by matching the frequency of the pre- and post-glitch solutions. Bottom: combined residuals from fits to independent timing models for the pre- and post-glitch intervals, respectively. The overall $\chi^{2}_{\nu}$ statistic for the fit is 2.20 for 32 DoF, taking into account the fit parameters for each interval.
Figure 4: Anti-glitch model for the spin-down of PSR J0821-4300. Top: pulse-phase residuals from the pre-glitch timing solution presented in Table 6. The glitch epoch of 2012 March 13 is estimated by matching the frequency of the pre- and post-glitch solutions. Bottom: combined residuals from fits to independent timing models for the pre- and post-glitch intervals, respectively. The overall $\chi^{2}_{\nu}$ statistic for the fit is 2.77 for 32 DoF, taking into account the fit parameters for each interval.
Figure 5: Pulse-phase residuals for PSR J0821-4300 using the alternative timing models presented in Table 7 that do not involve a glitch. Top: a quadratic model. Bottom: a cubic model (including the frequency second derivative).
2.3 PSR J1852++0040
PSR J1852++0040 was the first CCO to have its spin-down rate measured (Halpern & Gotthelf, 2010), but there followed a 12 yr hiatus in its monitoring. We have obtained a new set of observations of PSR J1852++0040 from a joint Chandra and XMM-Newton investigation designed to restart a phase-connected timing solution in 2021–2023 (Table 8). Photon arrival times were extracted in the 151-5 keV energy range using a $12^{\prime\prime}$ or $18^{\prime\prime}$ radius aperture with XMM-Newton, and a $1\farcs 8$ radius aperture with Chandra.
Table 9 presents the new results from the 10 observations obtained in 2021–2023, along with results from the 2004–2009 era (Halpern & Gotthelf, 2010). As shown in Figure 6, we are able to obtain independent timing solutions for PSR J1852++0040 both before and after the 12 yr gap. However, the long gap does not allow the cycle count to be maintained, given the uncertainty on the frequency derivative of the earlier solution.
During the 2021–2023 epoch, the frequency derivative is poorly constrained in the fit, as it contributes less than 2 cycles, based on the prior measured value in 2004–2009. The ephemeris for the earlier era predicts the frequency in 2021–2023 to within $\Delta f=4.05\times 10^{-9}$ Hz, comparable to the extrapolated uncertainty on this parameter ($\sigma_{f}=4.03\times\ 10^{-9}$ Hz). This is of a similar magnitude as that measured for both the 1E 1207.4-5209 and PSR J0821-4300 glitches. Thus, we would not be sensitive to a glitch in PSR J1852++0040 having similar magnitude as inferred for the two other CCO pulsars.
Table 8: Log of X-ray Timing Observations of PSR J1852++0040
Mission Instrument ObsID Date ExpaaExposure times for the XMM-Newton EPIC-pn do not reflect the 29% deadtime in SmallWindow (sw) mode.
/Mode (UT) (ks)
XMM EPIC-pn/sw 0204970201 2004 Oct 18 30.6
XMM EPIC-pn/sw 0204970301 2004 Oct 23 30.5
XMM EPIC-pn/sw 0400390201 2006 Oct 8 29.7
Chandra ACIS-S3/$C$ 6676 2006 Nov 23 32.2
XMM EPIC-pn/sw 0400390301 2007 Mar 20 30.5
Chandra ACIS-S3/$C$ 9101 2007 Nov 12 33.1
Chandra ACIS-S3/$C$ 9102 2008 Jun 16 31.2
XMM EPIC-pn/sw 0550670201 2008 Sep 19 21.2
XMM EPIC-pn/sw 0550670301 2008 Sep 21 31.0
XMM EPIC-pn/sw 0550670401 2008 Sep 23 34.8
XMM EPIC-pn/sw 0550670501 2008 Sep 29 33.0
XMM EPIC-pn/sw 0550670601 2008 Oct 10 36.0
Chandra ACIS-S3/$C$ 9823 2008 Nov 21 30.1
Chandra ACIS-S3/$C$ 9824 2009 Feb 20 29.6
XMM EPIC-pn/sw 0550671001 2009 Mar 16 27.0
XMM EPIC-pn/sw 0550670901 2009 Mar 17 26.0
XMM EPIC-pn/sw 0550671201 2009 Mar 23 27.3
XMM EPIC-pn/sw 0550671101 2009 Mar 25 19.9
XMM EPIC-pn/sw 0550671301 2009 Apr 4 26.0
XMM EPIC-pn/sw 0550671901 2009 Apr 10 30.5
XMM EPIC-pn/sw 0550671801 2009 Apr 22 28.0
Chandra ACIS-S3/$C$ 10128 2009 Jun 2 33.2
Chandra ACIS-S3/$C$ 10129 2009 Jul 29 32.2
New Observations – This Work
XMM EPIC-pn/sw 0872790101 2021 Mar 15 36.4
XMM EPIC-pn/sw 0872790201 2021 Mar 16 31.5
XMM EPIC-pn/sw 0872790301 2021 Mar 18 29.6
XMM EPIC-pn/sw 0872790401 2021 Mar 24 29.6
XMM EPIC-pn/sw 0872790501 2021 Apr 23 29.6
Chandra ACIS-S3/$C$ 23866 2021 Jul 8 21.0
Chandra ACIS-S3/$C$ 25085 2021 Jul 9 21.0
Chandra ACIS-S3/$C$ 23867 2022 May 12 43.0
Chandra ACIS-S3/$C$ 23868 2023 Apr 30 10.1
Chandra ACIS-S3/$C$ 27823 2023 May 2 31.8
Table 9: Timing Solutions for PSR J1852++0040footnotetext: Notes. Derived parameters ($B_{s}$, $\dot{E}$, $\tau_{c}$) are based on the 2004–2009 timing solution.
[…]
R.A. (J2000) $18^{\rm h}52^{\rm m}38^{\rm s}\!.57$
Decl. (J2000) $00^{\circ}40^{\prime}19^{\prime\prime}\!.8$
Surface dipole magnetic field, $B_{s}$ $3.1\times 10^{10}$ G
Spin-down luminosity, $\dot{E}$ $3.0\times 10^{32}$ erg s-1
Characteristic age, $\tau_{c}$ 192 Myr
Quadratic Timing Solution (2004–2009)
Epoch of ephemeris (MJD TDB) 54168.00000093
Span of ephemeris (MJD) 53296–55041
Frequency, $f$ 9.53174258208(17) s-1
Frequency derivative, $\dot{f}$ $7.882(81)\times 10^{-16}$ s-2
Period, $P$ 0.1049126108253(19) s
Period derivative, $\dot{P}$ $8.676(90)\times 10^{-18}$
$\chi^{2}_{\nu}[{\rm DoF}]$ 1.01[20]
Quadratic Timing Solution (2021–2023)
Epoch of ephemeris (MJD TDB) 59677.00000016
Span of ephemeris (MJD) 59288–60066
Frequency, $f$ 9.53174220286(34) s-1
Frequency derivative, $\dot{f}$ $8.62(45)\times 10^{-16}$ s-2
Period, $P$ 0.1049126149991(37) s
Period derivative, $\dot{P}$ $9.48(50)\times 10^{-18}$
$\chi^{2}_{\nu}[{\rm DoF}]$ 0.976[7]
3 Discussion
3.1 Do CCOs Glitch?
New timing observations reported here extend our previous ephemeris for 1E 1207.4-5209, (Gotthelf & Halpern, 2020) and reveal similar noisy timing behavior in PSR J0821-4300. However, because of the sparse sampling, and the relatively imprecise ToAs of thermal X-ray pulses compared with those of radio pulsars, these results do not conclusively prove that glitches occur in CCOs. A glitch is a priori surprising in a CCO, as discussed in (Gotthelf & Halpern, 2020), because CCOs have $|\dot{f}|<10^{-15}$ s-2. The glitch activity measured in pulsars is best correlated with $\dot{f}$, while it is rare that any pulsar with $|\dot{f}|<10^{-15}$ s-2 is seen to glitch. According to Antonopoulou et al. (2022), PSR B0410+69 is the only pulsar with $\dot{f}$ as small as $\approx-5\times 10^{-16}$ s-2 that was observed to glitch.
Ultimately, there is a limit in which a weak glitch can be either obscured or mimicked by timing noise, and that limit may have been reached in our study. Another manifestation of this ambiguity is the ability of an early anti-glitch to fit the timing of PSR J0821-4300 in our data almost as well as a later glitch does. We consider the anti-glitch fit for PSR J0821-4300 to be unconvincing as a physical model because it rests on a short pre-glitch span of data. While anti-glitches are sometimes seen in magnetars and accreting pulsars, they are rare in rotation-powered pulsars (Tuo et al., 2024; Panin & Sokolova, 2025). In Section 3.2 we discuss timing noise as an alternative to fits involving glitches.
With these caveats in mind, the implications of the changes in frequency derivative $\dot{f}$ required in our glitch fits deserve explicit mention. The newly extended post-glitch timing solution for 1E 1207.4-5209, by refining its frequency derivative from the previous study, revealed a large relative change of $\approx 9\%$ in $\dot{f}$ across the putative glitch. That is, $\Delta\dot{f}/\dot{f}\approx-0.09$, with $\dot{f}$ post-glitch becoming less negative. Note that the magnitude of $\Delta\dot{f}\sim 10^{-17}$ s-2 is itself not unprecedented across glitches (e.g., Figure 11 of Antonopoulou et al. 2022), but its relative change and sign are unusual. These are different from the more common glitches in which $\Delta\dot{f}/\dot{f}\approx 10^{-3}$ with $\dot{f}$ becoming more negative in 80% of the cases (Antonopoulou et al., 2022). While the change in $\dot{f}$ in 1E 1207.4-5209 is only a $4\sigma$ result, more definitive evidence of the same effect is found in the case of PSR J0821-4300; its $\dot{f}$ became $7\%$ less negative after the hypothesized glitch, the change being of $9\sigma$ significance.
If the surface dipole magnetic field strength $B_{s}$ is naively estimated as $\propto f^{-3/2}\dot{f}^{1/2}$, the fitted changes in $\dot{f}$ imply a decrease of $B_{s}$ by a few percent within a decade, a much shorter timescale than expected from the ages of the host SNRs. The sign of the change is also in conflict with a leading theory for the absence of old CCOs (Ho, 2011), which postulates that their weak external dipole field $B_{s}$ was caused by prompt burial of a normal $B$ field by supernova fallback debris, while reemergence (increase of $B_{s}$) by ohmic diffusion should follow on timescales of thousands of years (Muslimov & Page, 1995), allowing the CCOs to join the population of canonical pulsars in the $P-\dot{P}$ diagram.
It should be noted that comparable or larger variations in $\Delta\dot{f}/\dot{f}$ have been discovered from several pulsars in recent years, but they are all associated with changes in magnetospheric emission. These phenomena include 1) intermittent pulsars in which the radio pulsations turn off for periods of days to years (Kramer et al., 2006; Camilo et al., 2012; Lyne et al., 2017) while $|\dot{f}|$ decreases by $\geq 50\%$, 2) quasi-periodic changes in radio pulse profile that are associated with a few percent changes in $\dot{f}$ (Lyne et al., 2010), and 3) the radio quiet PSR J2021+4026 that has correlated states of $\gamma$-ray flux and $\dot{f}$, differing by $\approx 20\%$ and $\approx 10\%$, respectively (Allafort et al., 2013; Fiori et al., 2024). However, none of these effects are easily applied to CCOs, with their steady, surface thermal X-ray emission and pulse profiles, small spin-down power, and no apparent magnetospheric generation of particles and winds. More likely, therefore, the years-long variation of $\dot{f}$ in CCOs has its origin in internal properties that are in common with other young but more energetic pulsars, such as high temperature, strong intrinsic $B$ field and superfluid behavior; in some combination these properties may cause glitches (Ho, 2015) or timing noise.
3.2 Do CCOs have Unusual Timing Noise?
Figure 6: Pulse-phase residuals for PSR J1852++0040 from the quadratic spin-down model presented in Table 9, from two independent timing solutions corresponding to the intervals 2004–2009 (top) and 2021–2023 (bottom).
Fits to the spin-down of 1E 1207.4-5209 and PSR J0821-4300 using only $f$ and $\dot{f}$ without invoking glitches leave systematic trends in their phase residuals that may be regarded as timing noise. Here we attempt to evaluate how this compares to timing noise in canonical pulsars with similar spin-down properties. Several metrics of timing noise were reviewed by Namkham et al. (2019). The one most useful here is $\sigma_{\rm TN,2}$ from Shannon & Cordes (2010), in which
$\sigma_{\rm TN,2}^{2}=\sigma_{\rm R,2}^{2}-\sigma_{\rm W}^{2}.$ (1)
Here $\sigma_{\rm R,2}$ is the root-mean-square (rms) of the measured residuals from a quadratic fit, and $\sigma_{\rm W}$ is the typical uncertainty of a ToA. The subscripts R and W indicate a red or white noise process.
An advantage of this method is that it explicitly acknowledges the red noise character of timing noise by including in the analysis the total time span $T$ of the observations. Shannon & Cordes (2010) found for hundreds of canonical pulsars that the mean value of $\sigma_{\rm TN,2}$ scales with the spin parameters as
$\bar{\sigma}_{\rm TN,2}=C_{2}\,f^{\alpha}\,|\dot{f}|^{\beta}\,T^{\gamma}\ \ \mu{\rm s},$ (2)
where $C_{2}=41.7,\alpha=-0.9,\beta=1.00$, and $\gamma=1.9$. Recognizing that there is large scatter in $\sigma_{\rm TN,2}$ around a given mean $\bar{\sigma}_{\rm TN,2}$, Shannon & Cordes (2010) modeled its distribution as log-normal, and found that the standard deviation of ln($\sigma_{\rm TN,2}$) = 1.6.
[…]
Another characterization of timing noise simply uses the frequency second derivative (equivalent to the braking index) added to the fit as an approximate measure of the residuals. An example is the cubic fit of 1E 1207.4-5209 in Figure 2, which yields braking index $n=f\ddot{f}\dot{f}^{-2}=8.6\times 10^{6}$. In comparison, pulsars with positive $\ddot{f}$ generally have $n<2\times 10^{5}$ (Figure 5 of Namkham et al. 2019). In the case of PSR J0821-4300, it is interesting that fitting a second derivative of $\ddot{f}=1.0\times 10^{-25}$ s-3 barely improves the fit in Figure 5, while the equivalent braking index is $n=2.3\times 10^{5}$. For either pulsar, this method indicates large timing noise, but the meaning is not precise as it does not take into account the time span of the observations or the nature of the residuals remaining after the cubic fit. For completeness, we recall that an earlier parameterization of timing noise by Arzoumanian et al. (1994) used the frequency second derivative measured over a time span of $T=10^{8}$ s to define
$\Delta_{8}={\rm log}\left({1\over 6f}\,|\ddot{f}|\,T^{3}\right).$
However, we do not have enough precise ToAs within any $10^{8}$ s time span to measure a significant $\ddot{f}$.
Finally, we note that the fitted second derivatives representing timing noise are large enough, given the long time spans of the data, to slowly produce the large fractional changes in $\dot{f}$ that, in context of the glitch model, were constrained to be discrete events. Neither interpretation of the changes in $\dot{f}$ appears to be clearly preferable, although timing noise is perhaps simpler.
4 Conclusions
Continued timing of the known CCO pulsars establishes PSR J0821-4300 as the second one with timing noise or glitches, quantitatively similar to the behavior of 1E 1207.4-5209 established previously. We have taken care to ensure that all of the models applied to the timing residuals of a particular pulsar have the same set of cycle counts between observations. Thus, they are the same timing solution fitted with different analytic formulas, and not aliases. Even so, is difficult to distinguish glitches from timing noise in these CCOs because of the infrequent sampling, and the lower precision of the ToAs that can be obtained from these thermal X-ray pulsations compared to those of typical radio pulsars. Nevertheless, whichever description is adopted, the magnitude of the timing irregularities is greater than in rotation-powered pulsars with comparably small spin-down rate and as weak a dipole magnetic field strength as CCOs.
One quantitative outcome of the extended monitoring is that the spin-down rate $\dot{f}$ in 1E 1207.4-5209 and PSR J0821-4300 appears to slow by $7-9\%$ on a timescale of a decade. This is probably not a monotonic trend representing a decrease of the dipole $B$ field, but more likely a representative short-term fluctuation. In the absence of evidence for magnetospheric activity and intermittent mode changes such as are seen in certain variable rotation-powered pulsars, the origin of the timing variations in CCOS should be sought in the internal structure of these young NSs, which are likely to have $B$ fields much stronger than their external dipole component. Accretion torque noise from a residual fallback disk is even a possible contributor. Noisy timing behavior is an important feature of CCOs as a class, and it may inform theories of glitches and timing noise in the general pulsar population.
Support for this work was provided by NASA through XMM grants 80NSSC18K0452, 80NSSC19K0866, 80NSSC20K1310, 80NSSC21K0819, 80NSSC22K0867, NICER grants 80NSSC19K1461, 80NSSC21K0106, 80NSSC21K1884, 80NSSC22K1300, and Chandra Awards SAO GO7-18063X, SAO GO0-21059X, SAO GO1-22065X issued by the Chandra X-ray Observatory Center, which is operated by the Smithsonian Astrophysical Observatory for and on behalf of NASA under contract NAS8-03060. This investigation is based partly on observations obtained with XMM-Newton, an ESA science mission with instruments and contributions directly funded by ESA Member States and NASA. This research has made use of data and/or software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC.

### v2

@@
 Abstract
-We present a timing study of the three known central compact object (CCO) pulsars, isolated cooling neutron stars in supernova remnants, using Chandra, XMM-Newton and NICER observations spanning two decades.
-Relative to canonical young pulsars, CCOs are spinning down at a very slow rate $|\dot{f}|<10^{-15}$ s-2, implying a surface dipole magnetic field strength $B_{s}<10^{11}$ G that is too weak to account for their X-ray emitting hot spots.
+We present a timing study of the three known central compact object (CCO) pulsars, isolated cooling neutron stars in supernova remnants (SNRs), using observations from Chandra, XMM-Newton and NICER spanning two decades.
+Relative to canonical young pulsars, CCOs are spinning down at a very slow rate $|\dot{f}|<10^{-15}$ s-2, implying a surface dipole magnetic field strength $B_{s}<10^{11}$ G that is too weak to account for their X-ray-emitting hot spots.
 Two CCO pulsars with sufficiently long monitoring, 1E 1207.4-5209 and PSR J0821-4300, are seen to deviate from steady spin-down; their timing residuals can be modeled by one or more glitches in $f$ and $\dot{f}$, or alternatively by extreme timing noise.
@@
 Glitch activity and timing noise in large samples of rotation-powered pulsars correlate best with $\dot{f}$, while the timing irregularities of the first two CCOs are extreme compared to pulsars of the same $\dot{f}$.
-Nevertheless, timing activity in CCOs may arise from properties that they share with other young but more energetic pulsars: high internal temperature, strong buried magnetic field and superfluid behavior.
+Nevertheless, timing activity in CCOs may arise from properties that they share with other young but more energetic pulsars: high internal temperature, strong buried magnetic field, and superfluid behavior.
 Alternatively, continuing low-level accretion of supernova debris is not ruled out as a source of timing noise in CCOs.
 Unified Astronomy Thesaurus concepts: Pulsars (1306); Neutron stars (1108); Compact objects (288); Pulsar timing method (1305)
-1 Introduction
-Central compact objects (CCOs) comprise a dozen or so cooling neutron stars (NSs) in supernova remnants (SNRs), defined and distinguished from canonical, young rotation-powered pulsars by their steady surface thermal X-ray emission, lack of surrounding pulsar wind nebula, and non-detection at any other wavelength (Pavlov et al.
+I Introduction
+Central compact objects (CCOs), comprising a dozen or so cooling neutron stars (NSs) in supernova remnants (SNRs), are defined and distinguished from canonical, young rotation-powered pulsars by their steady surface thermal X-ray emission, lack of surrounding pulsar wind nebula, and nondetection at any other wavelength [^1] Recently, the first detection of weak radio pulsations from a CCO, 1E 1207.4-5209, was reported (Zhang et al., 2025).
+(Pavlov et al.
 2002; see De Luca 2017 for a review).
 Considering that their faintness and spectral softness are obstacles to widespread discovery in the Galactic plane, CCOs may be as common as more conspicuous NS classes, e.g., magnetars, and may represent a major channel of NS birth.
-Pulsations have been detected in only three of the CCOs: 1E 1207.4-5209 in SNR PKS 1209-51/52, PSR J0821-4300 in SNR Puppis A, and PSR J1852++0040 in SNR Kes 79.
-They have typical periods ($0.424,0.112,0.104$ s) but small spin-down rates $\dot{f}=(-1.2,-6.8,-7.9)\times 10^{-16}$ s-2, implying surface dipole magnetic field strengths of $(9.8,2.9,3.1)\times 10^{10}$ G, respectively, which are exceptionally weak for such young pulsars (Gotthelf et al., 2013; Halpern & Gotthelf, 2010).
-Furthermore, the surface magnetic field strength of 1E 1207.4-5209 inferred from spin-down, $B_{s}=9.8\times 10^{10}$ G, is close to $B\approx 8\times 10^{10}$ G, the value measured from its unique series of spectroscopic absorption features (Bignami et al., 2003), interpreted as the electron cyclotron fundamental at 0.7 keV and its harmonics.
-The similar values of field strength obtained by the two independent methods is convincing evidence of the weak surface $B$ field for at least this one CCO.
-The remainder of the CCOs, while having similar continuum spectral properties as the CCO pulsars, have eluded searches for pulsations (Alford & Halpern, 2023).
-It is postulated that they may have even weaker magnetic fields, more uniform surface temperature distributions, or an unfavorable viewing geometry.
-In fact, it is difficult to create the hot spots needed to produce the X-ray pulsations observed in CCOs assuming only their weak surface dipoles.
+X-ray pulsations, evidently due to nonuniform surface temperature, have been detected from only three of the CCOs: 1E 1207.4-5209 in SNR PKS 1209-51/52, PSR J0821-4300 in SNR Puppis A and PSR J1852++0040 in SNR Kes 79.
+They have typical spin periods $P=(0.424,0.112,0.104$) s but small spin-down rates $\dot{f}=(-1.2,-6.8,-7.9)\times 10^{-16}$ s-2, implying surface dipole magnetic field strengths of $(9.8,2.9,3.1)\times 10^{10}$ G, respectively, which are exceptionally weak for such young pulsars (Gotthelf et al., 2013; Halpern & Gotthelf, 2010).
+The ages of the pulsars are the same as those of their SNRs, estimated as $t=(14,10,5)\times 10^{3}$ yr for PKS 1209-51/52 (Eppens et al., 2024), Puppis A (Aruga et al., 2022) and Kes 79 (Mayer & Becker, 2021), respectively.
+In contrast, when a pulsar “characteristic age” is quoted, defined as $\tau_{c}\equiv P/2\dot{P}$, the value is in the range $(2-3)\times 10^{8}$ yr for CCOs.
+Because the relation between $\tau_{c}$ and true age $t$ is $t=\tau_{c}[1-(P_{0}/P)^{2}]$ for dipole braking with constant magnetic field, this spectacular difference simply indicates that the current spin period $P$ of the pulsar is likely identical to its otherwise unknown period at birth $P_{0}$ because its $B$ field is so weak.
+Crucially, the surface magnetic field strength of 1E 1207.4-5209 inferred from spin-down, $B_{s}=9.8\times 10^{10}$ G, is close to $B\approx 8\times 10^{10}$ G, the value measured from its unique series of spectroscopic absorption features (Bignami et al., 2003) interpreted as the electron cyclotron fundamental at 0.7 keV and its harmonics.
+That similar values of field strength are obtained by the two independent methods is persuasive evidence of the weak surface $B$ field for at least this CCO.
+The remainder of the CCOs, having similar continuum spectral properties, have eluded searches for X-ray pulsations (Alford & Halpern, 2023).
+It is postulated that they may have even weaker magnetic fields, more uniform surface temperature distributions, or unfavorable geometry of hot spots.
+Alford et al.
+(2022) derived that the hot spots on the surface of PSR J0821-4300 are probably close to its spin axis, and this could be the case for CCOs in general (Alford & Halpern, 2023), which would limit the possible amplitude and detectability of pulsation.
+A similar conclusion about the magnetic axis of 1E 1207.4-5209, that it is nearly aligned with the spin, was tentatively drawn from the polarization swing and width of its radio pulse (Zhang et al., 2025).
+It is even difficult to create the hot spots needed to produce X-ray pulsations in CCOs assuming only their weak surface dipoles.
 The only mechanism thought to be capable of creating a nonuniform surface temperature is anisotropic heat conduction in a strong magnetic field.
-As recounted in Gotthelf & Halpern (2018), models that invoke strong non-dipolar fields in the crust have long been studied.
-More recently, e.g., Igoshev et al.
+As recounted in Gotthelf & Halpern (2018), models that include strong nondipolar components, mainly toroidal ones in the crust, have long been studied.
+More recently, Igoshev et al.
 (2021) calculated the magnetothermal evolution of strong crustal fields generated by a stochastic dynamo, finding a tangled geometry with negligible dipole component.
-This creates factors of  2 difference in surface temperature on small scales, sufficient to produce the observed pulsations in CCOs.
-Postulating the existence of strong non-dipole crustal fields was also the key to an explanation for the conspicuous absence of CCO descendants, the large population of weak $B$-field pulsars that should remain in the same location of the $P-\dot{P}$ diagram after their SNRs have faded.
-If CCOs are actually born with a canonical NS magnetic dipole field that was promptly buried by fallback of a small amount $(\sim 10^{-4}\,M_{\odot})$ of supernova debris, the buried field will diffuse back to the surface on a timescale of $\sim 10^{5}$ years (Muslimov & Page, 1995; Bernal et al., 2010; Ho, 2011; Viganò & Pons, 2012).
+This creates factors of $\sim 2$ difference in surface temperature on small scales, sufficient to produce the observed pulsations in CCOs.
+Postulating a temporary suppression of the surface dipole field was the key to an explanation for the conspicuous absence of CCO descendants, the large population of weak $B$-field pulsars that should remain in the same location of the $P-\dot{P}$ diagram after their SNRs have faded.
+If CCOs are actually born with a canonical NS magnetic dipole field that was promptly buried by fallback of a small amount $(\sim 10^{-4}\,M_{\odot})$ of supernova debris, the buried field will diffuse back to the surface on a timescale of $\sim 10^{5}$ yr (Muslimov & Page, 1995; Bernal et al., 2010; Ho, 2011; Viganò & Pons, 2012).
 A CCO will move vertically up in the $P-\dot{P}$ diagram as its dipole field grows back (Bogdanov et al., 2014; Ho, 2015; Luo et al., 2015), merging with the population of ordinary radio pulsars.
 The discovery of a glitch and/or strong timing noise from 1E 1207.4-5209 (Gotthelf & Halpern, 2018, 2020) was of timely relevance to these theories of CCOs.
-Glitches are thought to result from either “starquakes”, stress relief of the NS crust, or from unpinning of superfluid vortices in the inner part of the NS crust (e.g., Link et al., 1998).
+Glitches are thought to result from “starquakes”, stress relief of the NS crust, or from unpinning of superfluid vortices in the inner part of the NS crust (e.g., Link et al., 1998).
 A review of glitch observations and theory is given by Antonopoulou et al.
-(2022), while a comprehensive compilation of glitch statistics can be found in Espinoza et al.
-(2011) and Fuentes et al.
+(2022), while comprehensive compilations of glitch statistics can be found in Espinoza et al.
+(2011), Fuentes et al.
+(2017) and Basu et al.
 (2022).
@@
 (2019).
-The timing irregularities in 1E 1207.4-5209 were extreme relative to the general pulsar population, in which glitch activity is correlated primarily with frequency derivative $\dot{f}$.
-Naively, glitches are unexpected in CCOs because of their small $\dot{f}$.
-However, the idea of stronger internal $B$ fields, needed to explain CCO surface hot spots and population statistics, may have an additional application in producing the timing irregularities of CCOs.
-Ho (2015) proposed that glitch activity could be triggered by the motion of strong magnetic fields through the NS crust, interacting with the neutron superfluid there, and that glitches would be a way of identifying CCO descendants.
+The timing irregularities in 1E 1207.4-5209 were extreme relative to the general pulsar population, in which glitch activity is mainly correlated with the frequency derivative $\dot{f}$.
+Glitches are therefore unexpected in CCOs due to their small $\dot{f}$.
+However, the idea of stronger internal $B$ fields, needed to explain the CCOs’ surface hot spots and population statistics, may find additional application in producing the timing irregularities of CCOs.
+Ho (2015) proposed that glitches could be triggered by the motion of strong magnetic fields through the NS crust, interacting with the neutron superfluid there, and that glitches would be a way of identifying CCO descendants.
 To add complexity, it has not been ruled out that an entirely different process, low-level accretion in the propeller regime, could cause the observed timing noise, specifically the observed changes in $\dot{f}$ (Halpern et al., 2007; Gotthelf & Halpern, 2018).
-This, because the required accretion rate of $\leq 10^{-11}$ g s-1 makes a negligible contribution to the observed luminosity, while the mass supply could be a residual fallback disk of $\sim 10^{-6}\,M_{\odot}$, much less mass than may have fallen back on the NS initially.
-The spin parameters of CCOs fall in regime where dipole braking and accretion torque noise could be comparable.
+This is because the required accretion rate of $\leq 10^{11}$ g s-1 makes a negligible contribution to the observed luminosity, while the mass supply could be a residual fallback disk of $\sim 10^{-6}\,M_{\odot}$, much less mass than may have fallen back on the NS initially.
+The spin parameters of CCOs fall into a regime where dipole braking and accretion torque noise could be comparable.
 The subject of this paper is the long-term spin-down evolution of all three CCO pulsars spanning the two decades since their discovery.
-In Section 2 we present new Chandra, XMM-Newton, and NICER timing observations which we use to update timing solutions, explore alternative timing models, and identify timing irregularites in PSR J0821-4300 for the first time.
-Section 3 discusses the implications of the candidate glitches and timing noise, and compares the CCO results with those of the general pulsar population.
-Section 4 summarizes our conclusions.
-2 Timing Analysis
+In Section II we present new observations from Chandra, XMM-Newton, and NICER that we use to update timing solutions, explore alternative timing models, and identify timing irregularities in PSR J0821-4300 for the first time.
+Section III discusses the implications of the candidate glitches and timing noise and compares the CCO results with those of the general pulsar population.
+Section IV summarizes our conclusions.
+II Timing Analysis
 We obtained new timing observations for the three CCO pulsars as part of our continued monitoring programs using the Chandra and XMM-Newton observatories.
@@
 With the onset of the NICER mission, we continued our coverage of 1E 1207.4-5209 using NICER exclusively.
-The previously published data sets also included here are fully described in our earlier work (1E 1207.4-5209 — Gotthelf & Halpern, 2007; Halpern & Gotthelf, 2011; Gotthelf et al., 2013; Halpern & Gotthelf, 2015; Gotthelf & Halpern, 2018, 2020), (PSR J0821-4300 — Gotthelf & Halpern, 2009; Gotthelf et al., 2013), and (PSR J1852++0040 — Gotthelf et al., 2005; Halpern et al., 2007; Halpern & Gotthelf, 2010).
-All data sets were reprocessed and reduced using the latest software for each mission, following the methods described in our earlier papers.
+The previously published datasets also included here are fully described in our previous work (1E 1207.4-5209 — Gotthelf & Halpern, 2007; Halpern & Gotthelf, 2011; Gotthelf et al., 2013; Halpern & Gotthelf, 2015; Gotthelf & Halpern, 2018, 2020), (PSR J0821-4300 — Gotthelf & Halpern, 2009; Gotthelf et al., 2013), and (PSR J1852++0040 — Gotthelf et al., 2005; Halpern et al., 2007; Halpern & Gotthelf, 2010).
+All of the observations used here are listed in the Appendix.
+All datasets were reprocessed and reduced using the latest software for each mission, following the methods described in our previous papers.
 Photon arrival times were converted to barycentric dynamical time (TDB) using the DE405 solar system ephemeris and the Chandra coordinates given in the tables below for each object.
@@
 Starting from the filtered event files, we measured the pulse period at each epoch using a periodogram search for the maximum power around the expected frequency, as informed by our earlier work.
-For 1E 1207.4-5209 and PSR J0821-4300, their pulse profiles are well-characterized by a sine function, and we computed phase zero from the Fourier coefficients of the unbinned photon arrival times.
-We define phase zero as the peak of the sine function (see Gotthelf & Halpern, 2020, for details).
+For 1E 1207.4-5209 and PSR J0821-4300, their pulse profiles are well characterized by a sinusoidal function, and we computed phase zero from the Fourier coefficients of the unbinned photon arrival times.
+For consistency with Gotthelf & Halpern (2020), we define phase zero as the minimum of the pulse for 1E 1207.4-5209.
+For PSR J0821-4300 we use the maximum of the pulse as phase zero.
 For PSR J1852++0040, which has a much broader pulse profile, a high-significance pulse profile template was generated by adding all the folded profiles, aligned by cross-correlating with an iterated template.
 We define phase zero for this CCO as the middle of the valley in the pulse profile.
-Utilizing the TEMPO software (Hobbs et al., 2006), we initiated a phase-connected timing solution to a subset of ToAs for each pulsar, using a model for the rotation phase of the pulsar that includes one or two of its frequency derivatives:
+Using the TEMPO software (Nice et al., 2015), we initiated a phase-connected timing solution to a subset of ToAs for each pulsar, using a model for the rotation phase of the pulsar that includes one or two of its frequency derivatives:
 $\phi(t)=\phi_{o}+f(t-t_{o})+{1\over 2}\dot{f}(t-t_{o})^{2}+{1\over 6}\ddot{f}(t-t_{o})^{3}.$
-Generally, we obtained optimally spaced sets of observations that allow us to start and to maintain a phase-connected timing solution that accounts for every turn of the pulsar, with increased accuracy, by bootstrapping to longer intervals.
-This is usually possible unless the interval to the next ToA is too large to predict its cycle count by less that $\sim 10\%$ in phase.
+Generally, we obtained optimally spaced sets of observations that allow us to start and maintain a phase-connected timing solution that accounts for every turn of the pulsar, with increased accuracy, by bootstrapping to longer intervals.
+This is usually possible unless the interval to the next ToA is too large to predict its cycle count by less than $\sim 10\%$ in phase.
 As described below, due to data gaps and timing irregularities, it was necessary at some point to restart a new solution for each pulsar to continue to follow and/or interpret its spin-down behavior.
 However, if the timing residuals begin to deviate significantly and systematically from the trial solution, it signifies that the ephemeris no longer predicts the pulsar’s spin-down, and more complex models need to be considered.
-2.1 1E 1207.4-5209
-In this work we use all available data sets for 1E 1207.4-5209 from XMM-Newton, Chandra, and NICER, as logged in Table 1.
+II.1 1E 1207.4-5209 in PKS 1209-51/52
+In this work, we use all available data sets for 1E 1207.4-5209 from XMM-Newton, Chandra, and NICER, as listed in the Appendix.
 These span 23 yr and include 3.5 yr of monitoring since our previously published results (Gotthelf & Halpern, 2020).
 The new observations used here were acquired exclusively using NICER.
-In the following, we extend the post-2015 glitch ephemeris as well as update the alternative models suggested by the earlier work.
+In the following, we extend the post-2015 glitch ephemeris and update the alternative models suggested by the earlier work.
 We do not repeat the previously reported fit for a possible pre-2015 glitch from 1E 1207.4-5209, as that result is unchanged.
-In the following analysis, we limit the photon energy range to 0.5–1.6 keV, optimal for the pulsar’s observed soft X-ray spectrum.
+In the following analysis, we limit the photon energy range to 0.5–1.6 keV, which is optimal for the pulsar’s observed soft X-ray spectrum.
+Figure 1 graphs the ToA phase residuals from the pre-glitch timing solution given in Table 1, developed using data points from 2002 to 2014 (as in Gotthelf & Halpern, 2020).
+As noted in our earlier work, the very first ToA, the Chandra observation of 2000, does not appear consistent with the preglitch timing solution and may indicate an even earlier glitch.
+We again excluded this data point from the fit but continue to plot it in Figure 1 for reference.
+After 2015, deviations from the expected phase of this solution were evident, and we were no longer able to extend the phase-connected solution forward.
+Instead, we postulated a glitch and found that the subsequent ToAs could be characterized by a new, postglitch phase-connected ephemeris.
+In this work, we have extended the postglitch ephemeris, which remains consistent with the earlier result and shows no evidence of a newer glitch.
+The updated change of frequency between the post- and preglitch timing solutions is $\Delta f=(3.87\pm 0.39)\times 10^{-9}$ Hz, with a glitch magnitude of $\Delta f/f_{\rm pred}=(1.64\pm 0.16)\times 10^{-9}$.
+The postglitch solution matches $f_{\rm pred}$, the frequency predicted by the preglitch solution on 2015 January 19 (MJD 57041), which we estimate as the epoch of the glitch (ignoring any short-time-scale recovery behavior, which would not have been sampled).
+Of note, the glitch magnitude for 1E 1207.4-5209 is typical of rotation-powered pulsars with strong magnetic fields and rapid spin-down, including the Crab.
+For reference, $\Delta f/f=3\times 10^{-9}$ is where the lower-amplitude peak in the bimodal distribution of glitch magnitudes is centered (Espinoza et al., 2011).
+The magnitude of the glitch is 23% smaller than the value quoted in Gotthelf & Halpern (2020).
+Nevertheless, this is the same postglitch timing solution as the one published previously in the sense that the cycle counts calculated over the common time span are the same.
+The fit statistic of the combined pre- and postglitch data sets is $\chi^{2}_{\nu}=1.49$ for 60 DoF.
+This should be compared to $\chi^{2}_{\nu}=1.44$ for 42 DoF previously reported for the shorter span.
+Figure 1: Glitch model for the spin-down of 1E 1207.4-5209.
+Top: pulse-phase residuals from the pre-glitch timing solution presented in Table 1.
+The glitch epoch of 2015 January 19 is estimated by matching the frequency of the pre- and postglitch solutions.
+The year 2000 Chandra data point is not included in the fit (see Section II.1 for details).
+Bottom: combined residuals from fits to independent timing models for the pre- and postglitch intervals, respectively.
+The overall $\chi^{2}_{\nu}$ statistic for the fit is 1.49 for 60 DoF, taking into account the fit parameters for each interval.
+Table 1: Glitch Ephemerides for 1E 1207.4-5209footnotetext: Notes. Derived parameters ($B_{s}$, $\dot{E}$, $\tau_{c}$) are based on the preglitch timing solution.
+[…]
+R.A. (J2000) $12^{\rm h}10^{\rm m}00^{\rm s}\!.91$
+Decl. (J2000) $52^{\circ}26^{\prime}28^{\prime\prime}\!.4$
+Surface dipole magnetic field, $B_{s}$ $9.8\times 10^{10}$ G
+Spin-down luminosity, $\dot{E}$ $1.1\times 10^{31}$ erg s-1
+Characteristic age, $\tau_{c}\equiv P/2\dot{P}$ 303 Myr
+Preglitch Timing Solution (2002–2014)
+Epoch of ephemeris (MJD TDB)bbEpoch of the ephemeris corresponds to the minimum of the pulse. 54547.00000198
+Span of ephemeris (MJD) 52266–56829
+Frequency, $f$ 2.357763492491(28) s-1
+Frequency derivative, $\dot{f}$ $1.2317(66)\times 10^{-16}$ s-2
+Period, $P$ 0.4241307506816(50) s
+Period derivative, $\dot{P}$ $2.216(12)\times 10^{-17}$
+$\chi^{2}_{\nu}[{\rm DoF}]$ 1.80[25]
+Postglitch Timing Solution (2016–2023)
+Epoch of ephemeris (MJD TDB)bbEpoch of the ephemeris corresponds to the minimum of the pulse. 58144.00000219
+Span of ephemeris (MJD) 57597–59986
+Frequency, $f$ 2.35776345898(14) s-1
+Frequency derivative, $\dot{f}$ $1.120(27)\times 10^{-16}$ s-2
+Period, $P$ 0.424130756710(26) s
+Period derivative, $\dot{P}$ $2.015(49)\times 10^{-17}$
+$\chi^{2}_{\nu}[{\rm DoF}]$ 1.26[35]
+Glitch epoch (MJD)ccEpoch of the glitch estimated by matching the frequency of the two timing solutions; this assumes a constant postglitch $\dot{f}$. 57041
+$\Delta f$ $3.87(39)\times 10^{-9}$ s-1
+$\Delta f/f_{\rm pred}$ $1.64(16)\times 10^{-9}$
+$\Delta\dot{f}$ $1.12(28)\times 10^{-17}$ s-2
+$\Delta\dot{f}/\dot{f}$ 0.091(23)-0.091(23)
+Figure 2: Pulse-phase residuals for 1E 1207.4-5209 using the alternative timing models presented in Table 2 that do not involve a glitch.
+These fits include the year 2000 Chandra data point (see Section II.1 for details).
+Top: A quadratic model fit leaves a sinusoidal-like oscillation in the phase residuals.
+Bottom: the residuals from a cubic model fit (including the frequency second derivative).
+Table 2: Alternative Timing Solutions for 1E 1207.4-5209
+[…]
+Quadratic Timing Solution (2000–2023)
+Epoch of ephemeris (MJD TDB)bbEpoch of the ephemeris corresponds to the minimum of the pulse. 55478.00000467
+Span of ephemeris (MJD) 51549–59986
+Frequency, $f$ 2.357763483778(12) s-1
+Frequency derivative, $\dot{f}$ $1.1128(15)\times 10^{-16}$ s-2
+Period, $P$ 0.4241307522490(21) s
+Period derivative, $\dot{P}$ $2.0018(28)\times 10^{-17}$
+$\chi^{2}_{\nu}[{\rm DoF}]$ 8.55[64]
+Cubic Timing Solution (2000–2023)
+Epoch of ephemeris (MJD TDB)bbEpoch of the ephemeris corresponds to the minimum of the pulse. 55478.00000453
+Span of ephemeris (MJD) 51549–59986
+Frequency, $f$ 2.357763483219(32) s-1
+Frequency derivative, $\dot{f}$ $1.1260(17)\times 10^{-16}$ s-2
+Frequency second derivative, $\ddot{f}$ $4.67(25)\times 10^{-26}$ s-3
+Period, $P$ 0.4241307523495(58) s
+Period derivative, $\dot{P}$ $2.0255(30)\times 10^{-17}$
+Period second derivative, $\ddot{P}$ $8.40(45)\times 10^{-27}$ s-1
+$\chi^{2}_{\nu}[{\rm DoF}]$ 3.17[63]
+Cubic Timing Solution (2002–2023)
+Epoch of ephemeris (MJD TDB)bbEpoch of the ephemeris corresponds to the minimum of the pulse. 55478.00000447
+Span of ephemeris (MJD) 52266–59986
+Frequency, $f$ 2.357763483142(35) s-1
+Frequency derivative, $\dot{f}$ $1.1318(20)\times 10^{-16}$ s-2
+Frequency second derivative, $\ddot{f}$ $5.51(29)\times 10^{-26}$ s-3
+Period, $P$ 0.42413075236(63) s
+Period derivative, $\dot{P}$ $2.0359(36)\times 10^{-17}$
+Period second derivative, $\ddot{P}$ $0.92(53)\times 10^{-26}$ s-1
+$\chi^{2}_{\nu}[{\rm DoF}]$ 2.72[62]
+The uncertainty on the frequency derivative of the revised postglitch timing solution has grown, perhaps an effect of timing noise.
+In addition, the postglitch frequency derivative is larger (less negative) than the preglitch value by $\approx 9\%$, albeit at the $4\sigma$ level.
+This change is much larger than the $\Delta\dot{f}/\dot{f}\sim 10^{-3}$ typically seen across glitches, which also suggests that timing noise may predominate.
+Moreover, the sign of the change of $\dot{f}$ is unusual and has implications that will be discussed in Section III.
+Weak glitches can be difficult to detect in the presence of timing noise, and timing noise can masquerade as glitches (Antonopoulou et al., 2022).
+A possible example of such an ambiguity is manifest in the downward curvature of the residuals around the time of the assumed glitch in Figure 1.
+Following Gotthelf & Halpern (2020), we therefore considered alternative timing models that do not require the use of a glitch to characterize the timing behavior of 1E 1207.4-5209.
+First, we show that a fit to the data using a quadratic model only, parameterized by the frequency and first derivative, and including the year 2000 observation, leaves a strong sinusoidal-like residual in the phases, with an amplitude of $\sim 0.1$ cycles (see Figure 2).
+As in Gotthelf & Halpern (2020), this results in an unacceptable fit statistic of $\chi^{2}_{\nu}=8.549$ for 64 DoF, substantially worse than for the glitch model.
+The sinusoidal residual with a period similar to the data span suggests that a higher-degree polynomial term should be added to the model.
+The lower panel of Figure 2 shows the result of fitting a cubic model that includes a frequency second derivative.
+This reduces the statistic to $\chi^{2}_{\nu}=3.17$ for 63 DoF, still significantly higher than for the glitch model.
+By excluding the 2000 data point, the fit statistic improves marginally to $\chi^{2}_{\nu}=2.72$ for 62 DoF but still suggests larger than expected timing noise, as will be discussed in Section III.
+The parameters of the quadratic and cubic fits are given in Table 2.
+It is important that each of these models, as well as the glitch model, have the same cycle count between all adjacent ToAs.
+Thus, they are all the same phase-connected solution that we believe is the true one, not an alias.
+Finally, we note that the binary orbit model suggested in Gotthelf & Halpern (2020) remains a viable interpretation, although less likely a priori.
+II.2 PSR J0821-4300 in Puppis A
+Figure 3: Glitch model for the spin-down of PSR J0821-4300.
+Top: pulse-phase residuals from the preglitch timing solution presented in Table 3.
+The glitch epoch of 2015 December 6 is estimated by matching the frequency of the pre- and postglitch solutions.
+Bottom: combined residuals from fits to independent timing models for the pre- and postglitch intervals, respectively.
+The overall $\chi^{2}_{\nu}$ statistic for the fit is 2.20 for 32 DoF, taking into account the fit parameters for each interval.
+Table 3: Glitch Ephemerides for PSR J0821-4300footnotetext: Notes.
+Position and proper motion from Mayer et al.
+(2020); Derived parameters ($B_{s}$, $\dot{E}$, $\tau_{c}$) are based on the preglitch timing solution.
+[…]
+Epoch of position and $\mu$ (MJD) 53964.0
+R.A. (J2000) $08^{\rm h}21^{\rm m}57.\!^{\rm s}3653(31)$
+Decl. (J2000) $43^{\circ}00^{\prime}17.\!^{\prime\prime}074(43)$
+R.A. proper motion, $\mu_{\alpha}\,{\rm cos}\,\delta$ $54.1\pm 8.3$ mas yr-1
+Decl. proper motion, $\mu_{\delta}$ $28.1\pm 10.5$ mas yr-1
+Surface dipole magnetic field, $B_{s}$ $2.9\times 10^{10}$ G
+Spin-down luminosity, $\dot{E}$ $1.9\times 10^{32}$ erg s-1
+Characteristic age, $\tau_{c}$ 254 Myr
+Preglitch Timing Solution (2010–2014)
+Epoch of ephemeris (MJD TDB)bbEpoch of the ephemeris corresponds to the maximum of the pulse. 56065.00000030
+Span of ephemeris (MJD) 55182–56948
+Frequency, $f$ 8.865291025552(85) s-1
+Frequency derivative, $\dot{f}$ $6.801(44)\times 10^{-16}$ s-2
+Period, $P$ 0.1127994554401(11) s
+Period derivative, $\dot{P}$ $8.654(56)\times 10^{-18}$
+$\chi^{2}_{\nu}[{\rm DoF}]$ 2.70[19]
+Postglitch Timing Solution (2015–2022)
+Epoch of ephemeris (MJD TDB)bbEpoch of the ephemeris corresponds to the maximum of the pulse. 58530.00000022
+Span of ephemeris (MJD) 57272–59787
+Frequency, $f$ 8.865290891814(75) s-1
+Frequency derivative, $\dot{f}$ $6.321(27)\times 10^{-16}$ s-2
+Period, $P$ 0.11279945714171(95) s
+Period derivative, $\dot{P}$ $8.042(34)\times 10^{-18}$
+$\chi^{2}_{\nu}[{\rm DoF}]$ 1.47[13]
+Glitch epoch (MJD)ccEpoch of the glitch estimated by matching the frequency of the two timing solutions; this assumes a constant postglitch $\dot{f}$. 57362
+$\Delta f$ $6.26(55)\times 10^{-9}$ s-1
+$\Delta f/f_{\rm pred}$ $7.06(62)\times 10^{-10}$
+$\Delta\dot{f}$ $4.80(52)\times 10^{-17}$ s-2
+$\Delta\dot{f}/\dot{f}$ 0.0706(76)-0.0706(76)
+Figure 4: Antiglitch model for the spin-down of PSR J0821-4300.
+Top: pulse-phase residuals from the preglitch timing solution presented in Table 4.
+The glitch epoch of 2011 September 14 is estimated by matching the frequency of the pre- and postglitch solutions.
+Bottom: combined residuals from fits to independent timing models for the pre- and postglitch intervals, respectively.
+The overall $\chi^{2}_{\nu}$ statistic for the fit is 2.77 for 32 DoF, taking into account the fit parameters for each interval.
+Table 4: Antiglitch Ephemerides for PSR J0821-4300
+[…]
+Preglitch Timing Solution (2010–2011)
+Epoch of ephemeris (MJD TDB)bbEpoch of the ephemeris corresponds to the maximum of the pulse. 55484.00000014
+Span of ephemeris (MJD) 55182–55784
+Frequency, $f$ 8.86529106058(42) s-1
+Frequency derivative, $\dot{f}$ $6.14(50)\times 10^{-16}$ s-2
+Period, $P$ 0.1127994549943(53) s
+Period derivative, $\dot{P}$ $7.81(64)\times 10^{-18}$
+$\chi^{2}_{\nu}[{\rm DoF}]$ 2.535[11]
+Postglitch Timing Solution (2012–2022)
+Epoch of ephemeris (MJD TDB)bbEpoch of the ephemeris corresponds to the maximum of the pulse. 57830.00000060
+Span of ephemeris (MJD) 55873–59787
+Frequency, $f$ 8.865290929387(33) s-1
+Frequency derivative, $\dot{f}$ $6.2179(81)\times 10^{-16}$ s-2
+Period, $P$ 0.11279945666364(42) s
+Period derivative, $\dot{P}$ $7.912(10)\times 10^{-18}$
+$\chi^{2}_{\nu}[{\rm DoF}]$ 2.695[21]
+Glitch epoch (MJD)ccEpoch of the glitch estimated by matching the frequency of the two timing solutions; this assumes a constant postglitch $\dot{f}$. 55828
+$\Delta f$ $5.4(6)\times 10^{-9}$ s-1
+$\Delta f/f_{\rm pred}$ $6.1(6)\times 10^{-10}$
+As listed in the Appendix, the timing observations of PSR J0821-4300 used in this work include the Chandra and XMM-Newton data sets previously reported in Gotthelf et al.
+(2013) that span the years 2009–2012, but exclude the two observations from 2001.
+In retrospect, it is not possible to span the large time gap from 2001 to 2009 due to the uncertainty in extrapolation of the preglitch ephemeris.
+In creating ToAs for PSR J0821-4300, we restricted the extracted photon energies to the 1.5–4.5 keV range because of the unique energy dependence of the pulse phase.
+Below 1.5 keV, the pulse is shifted $\approx 180^{\circ}$, effectively canceling the pulsations if the full energy band is used (see Gotthelf & Halpern, 2009, for details).
+Photon arrival times were extracted using circular apertures of radius $30^{\prime\prime}$ and $1\farcs 2$ for the XMM-Newton and Chandra datasets, respectively.
+Following the results presented in Gotthelf et al.
+(2013), we were able to continue monitoring this pulsar and extend its ephemeris but only up to 2014.
+Subsequently, the yearly monitoring showed increasing departures of the pulse arrival times from the refined ephemeris.
+Motivated by the similarity to the glitch-like behavior of 1E 1207.4-5209, we then initiated a renewed joint XMM-Newton and Chandra timing campaign on PSR J0821-4300 in 2019-2020 to generate an independent phase-connected timing solution for comparison.
+In Figure 3, we postulate that a glitch in pulse arrival time likely occurred between the 2015 and 2016 observations.
+Indeed, the monotonic deviations from the preglitch timing solution given in Table 3 imply a change in frequency of magnitude $\Delta f/f_{\rm pred}=7.1(6)\times 10^{-10}$, a value similar to that found for 1E 1207.4-5209.
+Including the pre- and postglitch timing solutions, the statistic for the combined data set is $\chi^{2}_{\nu}=2.20$ for 32 DoF, taking into account the fit parameters for each interval.
+The sparse coverage around the time of the glitch, between 2014 and 2017, prevents a definitive description of the glitch behavior.
+However, formal fits show curvature in the trend of postglitch residuals relative to the preglitch solution, corresponding to a $\approx 7\%$ change in frequency derivative, which becomes less negative, similar to the case of 1E 1207.4-5209 but at a higher $9\sigma$ level of significance (see Table 3).
+Figure 5: Pulse-phase residuals for PSR J0821-4300 using the alternative timing models presented in Table 5 that do not involve a glitch.
+Top: a quadratic model.
+Bottom: a cubic model (including the frequency second derivative).
+Table 5: Alternative Timing Solutions for PSR J0821-4300
+[…]
+Quadratic Timing Solution (2010–2022)
+Epoch of ephemeris (MJD TDB)bbEpoch of the ephemeris corresponds to the maximum of the pulse. 57485.00000096
+Span of ephemeris (MJD) 55182–59787
+Frequency, $f$ 8.865290948152(24) s-1
+Frequency derivative, $\dot{f}$ $6.2723(51)\times 10^{-16}$ s-2
+Period, $P$ 0.11279945642488(30) s
+Period derivative, $\dot{P}$ $7.9807(65)\times 10^{-18}$
+$\chi^{2}_{\nu}[{\rm DoF}]$ 6.79[35]
+Cubic Timing Solution (2010–2022)
+Epoch of ephemeris (MJD TDB)bbEpoch of the ephemeris corresponds to the maximum of the pulse. 57485.00000097
+Span of ephemeris (MJD) 55182–59787
+Frequency, $f$ 8.865290947727(71) s-1
+Frequency derivative, $\dot{f}$ $6.2665(52)\times 10^{-16}$ s-2
+Frequency second derivative, $\ddot{f}$ $1.00(16)\times 10^{-25}$ s-3
+Period, $P$ 0.11279945643030(91) s
+Period derivative, $\dot{P}$ $7.9733(66)\times 10^{-18}$
+Period second derivative, $\ddot{P}$ $1.27(20)\times 10^{-27}$ s-1
+$\chi^{2}_{\nu}[{\rm DoF}]$ 5.81[34]
+Curiously, a model involving an antiglitch (decrease in frequency) fits nearly as well as the glitch model, with the event occurring at an earlier time, around 2011 September (Figure 4).
+The parameters of this model are given in Table 4, where it can be seen that the magnitude of the antiglitch is similar to that of the glitch model, while the statistic for the combined data set is slightly worse at $\chi^{2}_{\nu}=2.77$.
+Since the preglitch segment in this model is very short, the initial frequency derivative was only measured to a precision of $\approx 8\%$, but it is consistent with the postglitch value.
+Note that this model has the same set of cycle counts as the glitch model, despite being a different analytic and qualitative description of the same timing data.
+As in the case of 1E 1207.4-5209, the smooth curvature of the residuals around the epoch of the assumed glitch in Figure 3 leads us to consider alternative timing models without the use of a glitch to characterize the timing of PSR J0821-4300.
+A basic quadratic model leaves phase residuals with $\sim 0.1$ cycle amplitude, similar in magnitude to those of 1E 1207.4-5209, but not clearly sinusoidal in nature (Figure 5).
+The fit statistic is $\chi^{2}_{\nu}=6.79$ for 35 DoF, again substantially worse than for the glitch model and similar to that found for 1E 1207.4-5209.
+The addition of a second frequency derivative term for a cubic model has little effect on the shape of the residual curve, and only reduces $\chi^{2}_{\nu}$ to 5.80 for 34 DoF.
+Parameters from the quadratic and cubic fits are given in Table 5.
+It can be seen that the residual curve in Figure 5 could also be interpreted as an early antiglitch or a later glitch.
+Again, none of these are distinct timing solutions (aliases); they all have the same set of cycle counts.
+II.3 PSR J1852++0040 in Kes 79
+PSR J1852++0040 was the first CCO to have its spin-down rate measured (Halpern & Gotthelf, 2010), but there followed a 12 yr hiatus in its monitoring.
+As listed in the Appendix, we obtained a new set of observations of PSR J1852++0040 from a joint Chandra and XMM-Newton investigation designed to restart a phase-connected timing solution in 2021–2023.
+Photon arrival times were extracted in the 1–5 keV energy range using a $12^{\prime\prime}$ or $18^{\prime\prime}$ radius aperture with XMM-Newton, and a $1\farcs 8$ radius aperture with Chandra.
+Table 6 presents the new results from the 10 observations obtained in 2021–2023, along with results from the 2004–2009 era (Halpern & Gotthelf, 2010).
+As shown in Figure 6, we are able to obtain independent timing solutions for PSR J1852++0040 both before and after the 12 yr gap.
+However, the long gap does not allow the cycle count to be maintained, given the uncertainty on the frequency derivative of the earlier solution.
+During the 2021–2023 epoch, the frequency derivative is poorly constrained in the fit, as it contributes less than 2 cycles, based on the prior measured value in 2004–2009.
+The ephemeris for the earlier era predicts the frequency in 2021–2023 to within $\Delta f=4.05\times 10^{-9}$ Hz, comparable to the extrapolated uncertainty on this parameter ($\sigma_{f}=4.03\times\ 10^{-9}$ Hz).
+This is of a similar magnitude as that measured for both the 1E 1207.4-5209 and PSR J0821-4300 glitches.
+Thus, we would not be sensitive to a glitch in PSR J1852++0040 having similar magnitude as inferred for the two other CCO pulsars.
+Figure 6: Pulse-phase residuals for PSR J1852++0040 from the quadratic spin-down model presented in Table 6, from two independent timing solutions corresponding to the intervals 2004–2009 (top) and 2021–2023 (bottom).
+Table 6: Timing Solutions for PSR J1852++0040footnotetext: Notes. Derived parameters ($B_{s}$, $\dot{E}$, $\tau_{c}$) are based on the 2004–2009 timing solution.
+[…]
+R.A. (J2000) $18^{\rm h}52^{\rm m}38^{\rm s}\!.57$
+Decl. (J2000) $00^{\circ}40^{\prime}19^{\prime\prime}\!.8$
+Surface dipole magnetic field, $B_{s}$ $3.1\times 10^{10}$ G
+Spin-down luminosity, $\dot{E}$ $3.0\times 10^{32}$ erg s-1
+Characteristic age, $\tau_{c}$ 192 Myr
+Quadratic Timing Solution (2004–2009)
+Epoch of ephemeris (MJD TDB)bbEpoch of the ephemeris corresponds to the minimum of the pulse. 54168.00000093
+Span of ephemeris (MJD) 53296–55041
+Frequency, $f$ 9.53174258208(17) s-1
+Frequency derivative, $\dot{f}$ $7.882(81)\times 10^{-16}$ s-2
+Period, $P$ 0.1049126108253(19) s
+Period derivative, $\dot{P}$ $8.676(90)\times 10^{-18}$
+$\chi^{2}_{\nu}[{\rm DoF}]$ 1.01[20]
+Quadratic Timing Solution (2021–2023)
+Epoch of ephemeris (MJD TDB)bbEpoch of the ephemeris corresponds to the minimum of the pulse. 59677.00000016
+Span of ephemeris (MJD) 59288–60066
+Frequency, $f$ 9.53174220286(34) s-1
+Frequency derivative, $\dot{f}$ $8.62(45)\times 10^{-16}$ s-2
+Period, $P$ 0.1049126149991(37) s
+Period derivative, $\dot{P}$ $9.48(50)\times 10^{-18}$
+$\chi^{2}_{\nu}[{\rm DoF}]$ 0.976[7]
+III Discussion
+III.1 Do CCOs Glitch?
+The new timing observations reported here extend our previous ephemeris for 1E 1207.4-5209, (Gotthelf & Halpern, 2020) and reveal a similar noisy timing behavior in PSR J0821-4300.
+However, because of the sparse sampling and the relatively imprecise ToAs of thermal X-ray pulses compared with those of radio pulsars, these results do not conclusively prove that glitches occur in CCOs.
+A glitch is a priori surprising in a CCO, as discussed in Gotthelf & Halpern (2020), because CCOs have $|\dot{f}|<10^{-15}$ s-2.
+The glitch activity parameter for an individual pulsar is defined as
+$\dot{f}_{g}\equiv{\sum_{j}\Delta f_{j}\over T},$
+where the numerator is the sum of the frequency steps $\Delta f_{j}$ over glitches and $T$ is the total time span of monitoring; therefore, $\dot{f}_{g}/|\dot{f}|$ approximates the long-term fraction of the spin-down that is reversed by glitches.
+Basu et al.
+(2022) find that $\dot{f}_{g}$ correlates best with $\dot{f}$ (and inversely with $\tau_{c}$) such that $\dot{f}_{g}/|\dot{f}|=0.018\pm 0.003$ in the range $10^{-14}<|\dot{f}|<10^{-10.5}$ s-2.
+However, it is rare that any pulsar with $|\dot{f}|<10^{-15}$ s-2 will glitch.
+As summarized in Basu et al.
+(2022), the proportionality of glitch activity to spin-down rate can be understood as reflecting the differential velocity that must build up between superfluid and normal components before stress becomes large enough to trigger the unpinning of superfluid vortices.
+The same could apply if unpinning is triggered by a crustquake.
+According to Antonopoulou et al.
+(2022), PSR B0410+69 is the only pulsar with $\dot{f}$ as small as $\approx-5\times 10^{-16}$ s-2 that was observed to glitch.
+CCOs, which have a similar or smaller spin-down rate, are statistically not expected to glitch.
+Ultimately, there is a limit in which a weak glitch can be either obscured or mimicked by timing noise, and that limit may have been reached in our study.
+Another manifestation of this ambiguity is the ability of an early antiglitch to fit the timing of PSR J0821-4300 in our data almost as well as a later glitch does.
+We consider the antiglitch fit for PSR J0821-4300 to be unconvincing as a physical model because it rests on a short preglitch span of data.
+Although antiglitches are sometimes seen in magnetars and accreting pulsars, they are rare in rotation-powered pulsars (PSR B0540-69: Tuo et al.
+2024, PSR J1522-5735: Panin & Sokolova 2025).
+Also, Espinoza et al.
+(2024) favor a fit of the same event in PSR B0540-69 as changes in $\dot{f}$ with no evident step in $f$, i.e., similar to timing noise.
+In Section III.2 we discuss timing noise as an alternative to fits involving glitches.
+With these caveats in mind, the implications of the changes in frequency derivative $\dot{f}$ required in our glitch fits deserve explicit mention.
+The newly extended postglitch timing solution for 1E 1207.4-5209, by refining its frequency derivative from the previous study, revealed a large relative change of $\approx 9\%$ in $\dot{f}$ across the putative glitch.
+That is, $\Delta\dot{f}/\dot{f}\approx-0.09$, with $\dot{f}$ postglitch becoming less negative.
+Note that the magnitude of $\Delta\dot{f}\sim 10^{-17}$ s-2 is itself not unprecedented across glitches (e.g., Figure 11 of Antonopoulou et al.
+2022), but its relative change and sign are unusual.
+These are different from the more common glitches in which $\Delta\dot{f}/\dot{f}\approx 10^{-3}$ with $\dot{f}$ becoming more negative in 80% of the cases (Antonopoulou et al., 2022).
+Although the change in $\dot{f}$ in 1E 1207.4-5209 is only a $4\sigma$ result, more definitive evidence of the same effect is found in the case of PSR J0821-4300; its $\dot{f}$ became $7\%$ less negative after the hypothesized glitch, the change being of $9\sigma$ significance.
+If the surface dipole magnetic field strength $B_{s}$ is naively estimated as $\propto f^{-3/2}\dot{f}^{1/2}$, the fitted changes in $\dot{f}$ imply a decrease of $B_{s}$ by a few percent within a decade, a much shorter timescale than expected from the ages of the host SNRs.
+The sign of the change is also in conflict with a leading theory for the absence of old CCOs (Ho, 2011), which postulates that their weak external dipole field $B_{s}$ was caused by prompt burial of a normal $B$ field by supernova fallback debris, while reemergence (increase of $B_{s}$) by ohmic diffusion should follow on timescales of thousands of years (Muslimov & Page, 1995), allowing the CCOs to join the population of canonical pulsars in the $P-\dot{P}$ diagram.
+It should be noted that comparable or larger variations in $\Delta\dot{f}/\dot{f}$ have been discovered from several pulsars in recent years, but they are all associated with changes in magnetospheric emission.
+These phenomena include (1) intermittent pulsars in which the radio pulsations turn off for periods of days to years (Kramer et al., 2006; Camilo et al., 2012; Lyne et al., 2017) while $|\dot{f}|$ decreases by $\geq 50\%$, (2) quasiperiodic changes in radio pulse profile that are associated with a few percent changes in $\dot{f}$ (Lyne et al., 2010), and (3) the radio-quiet PSR J2021+4026 that has correlated states of $\gamma$-ray flux and $\dot{f}$, differing by $\approx 20\%$ and $\approx 10\%$, respectively (Allafort et al., 2013; Fiori et al., 2024).
+However, none of these effects can easily be applied to CCOs, with their steady surface thermal X-ray emission and pulse profiles, small spin-down power, and no apparent magnetospheric generation of particles and winds.
+More likely, therefore, the years-long variation of $\dot{f}$ in CCOs has its origin in internal properties that are in common with other young but more energetic pulsars, such as high temperature, strong intrinsic $B$ field and superfluid behavior; in some combination, these properties may cause glitches (Ho, 2015) or timing noise.
+However, with the discovery of weak radio pulsations from 1E 1207.4-5209 (Zhang et al., 2025), especially if they are intermittent, magnetospheric effects may be reconsidered as a driver of variations in $\dot{f}$.
+III.2 Do CCOs have Unusual Timing Noise?
+Fits to the spin-down of 1E 1207.4-5209 and PSR J0821-4300 using only $f$ and $\dot{f}$ without invoking glitches leave systematic trends in their phase residuals that may be regarded as timing noise.
+Here we attempt to evaluate how this compares to timing noise in canonical pulsars with similar spin-down properties.
+Several metrics of timing noise were reviewed by Namkham et al.
+(2019).
+The most useful here is $\sigma_{\rm TN,2}$ from Shannon & Cordes (2010), in which
+$\sigma_{\rm TN,2}^{2}=\sigma_{\rm R,2}^{2}-\sigma_{\rm W}^{2}.$ (1)
+Here $\sigma_{\rm R,2}$ is the rms of the residuals measured from a quadratic fit, and $\sigma_{\rm W}$ is the typical uncertainty of a ToA.
+The subscripts R and W indicate a red or white noise process.
+An advantage of this method is that it explicitly acknowledges the red noise character of the timing noise by including in the analysis the total monitoring time span $T$.
+Shannon & Cordes (2010) found for hundreds of canonical pulsars that the mean value of $\sigma_{\rm TN,2}$ scales with the spin parameters as
+$\bar{\sigma}_{\rm TN,2}=C_{2}\,f^{\alpha}\,|\dot{f}|^{\beta}\,T^{\gamma}\ \ \mu{\rm s},$ (2)
+where $C_{2}=41.7,\alpha=-0.9,\beta=1.0$, and $\gamma=1.9$.
+Recognizing the large scatter in $\sigma_{\rm TN,2}$ around a given mean $\bar{\sigma}_{\rm TN,2}$, Shannon & Cordes (2010) modeled its distribution as log-normal and found that the standard deviation of ln($\sigma_{\rm TN,2}$) = 1.6.
+[…]
+[…]
+Another characterization of timing noise simply uses the frequency second derivative (equivalent to the braking index) added to the fit as an approximate measure of the residuals.
+An example is the cubic fit of 1E 1207.4-5209 in Figure 2, which yields a braking index of $n=f\ddot{f}\dot{f}^{-2}=8.6\times 10^{6}$.
+In comparison, pulsars with positive $\ddot{f}$ generally have $n<2\times 10^{5}$ (Figure 5 of Namkham et al.
+2019).
+In the case of PSR J0821-4300, it is interesting that fitting a second derivative of $\ddot{f}=1.0\times 10^{-25}$ s-3 barely improves the fit in Figure 5, while the equivalent braking index is $n=2.3\times 10^{5}$.
+For either pulsar, this method indicates large timing noise, but the meaning is not precise as it does not take into account the time span of the observations or the nature of the residuals remaining after the cubic fit.
+For completeness, we recall that an earlier parameterization of timing noise by Arzoumanian et al.
+(1994) used the frequency second derivative measured over a time span of $T=10^{8}$ s to define
+$\Delta_{8}={\rm log}\left({1\over 6f}\,|\ddot{f}|\,T^{3}\right).$
+However, we do not have enough precise ToAs within any time span of $10^{8}$ s to measure a significant $\ddot{f}$.
+Finally, we note that the fitted second derivatives representing timing noise are large enough, given the long time spans of the data, to slowly produce the large fractional changes in $\dot{f}$ that, in context of the glitch model, were constrained to be discrete events.
+Neither interpretation of the changes in $\dot{f}$ appears to be clearly preferable, although timing noise is perhaps simpler.
+IV Conclusions
+Continued timing of the known CCO pulsars establishes PSR J0821-4300 as the second one with timing noise or glitches, quantitatively similar to the behavior of 1E 1207.4-5209 established previously.
+We have taken care to ensure that all of the models applied to the timing residuals of a particular pulsar have the same set of cycle counts between observations.
+Thus, they are the same timing solution fitted with different analytic formulas and not aliases.
+Even so, distinguishing glitches from timing noise in these CCOs is difficult because of the infrequent sampling and lower precision of the ToAs that can be obtained from these thermal X-ray pulsations compared to those of typical radio pulsars.
+Nevertheless, whichever description is adopted, the magnitude of the timing irregularities is greater than in rotation-powered pulsars with a comparably small spin-down rate and a dipole magnetic field strength as small as that of CCOs.
+A quantitative result of extended monitoring is that the spin-down rate $\dot{f}$ in 1E 1207.4-5209 and PSR J0821-4300 appears to decrease by $7\%-9\%$ on a timescale of a decade.
+This is probably not a monotonic trend that represents a decrease in the dipole $B$ field but more likely a representative short-term fluctuation.
+In the absence of evidence for magnetospheric activity and intermittent mode changes such as are seen in certain variable rotation-powered pulsars, the origin of the timing variations in CCOs should be sought in the internal structure of these young NSs, which may have superfluid dynamics affected by high temperatures and $B$ fields much stronger than their external dipole component.
+Accretion torque noise from a residual fallback disk is even a possible contributor.
+Noisy timing behavior is an important feature of CCOs as a class, and it may inform theories of glitches and timing noise in the general pulsar population.
+This paper employs a list of Chandra datasets, obtained by the Chandra X-ray Observatory, contained in the Chandra Data Collection  https://doi.org/10.25574/cdc.523 (catalog DOI: 10.25574/cdc.523).
+Support for this work was provided by NASA through XMM grants 80NSSC18K0452, 80NSSC19K0866, 80NSSC20K1310, 80NSSC21K0819, 80NSSC22K0867; NICER grants 80NSSC19K1461, 80NSSC21K0106, 80NSSC21K1884, 80NSSC22K1300; and Chandra Awards SAO GO7-18063X, SAO GO0-21059X, SAO GO1-22065X issued by the Chandra X-ray Observatory Center, which is operated by the Smithsonian Astrophysical Observatory for and on behalf of NASA under contract NAS8-03060.
+This investigation is based partly on observations obtained with XMM-Newton, an ESA science mission with instruments and contributions directly funded by ESA Member States and NASA.
+This research has made use of data and/or software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC.
+The logs of observations of 1E 1207.4-5209, PSR J0821-4300, and PSR J1852++0040, are presented in Tables 78 and 9, respectively.
 […]
 Table 7: Log of X-ray Timing Observations of 1E 1207.4-5209
-Mission Instrument ObsIDaaThe NICER data sets are denoted by the ObsID and date of the
-first of the concatenated set of short adjacent observations.
-Date ExpbbExposure times for the XMM-Newton EPIC-pn do not reflect the
-29% deadtime in SmallWindow (sw) mode.
-/Mode (UT) (ks)
+Mission Instrument/Mode ObsIDaaThe NICER datasets are denoted by the ObsID and date of the first of the concatenated set of short adjacent observations.
+Date ExpbbExposure times for the XMM-Newton EPIC-pn do not reflect the 29% deadtime in SmallWindow (sw) mode.
+(UT) (ks)
 Chandra ACIS-S3/$C$ 751 2000 Jan 6 32.4
@@
 NICER XTI 4614011201-03 2023 Feb 11 15.8
-Table 2: Glitch Ephemerides for 1E 1207.4-5209footnotetext: Notes. Derived parameters ($B_{s}$, $\dot{E}$, $\tau_{c}$) are based on the pre-glitch timing solution.
-[…]
-R.A. (J2000) $12^{\rm h}10^{\rm m}00^{\rm s}\!.91$
-Decl. (J2000) $52^{\circ}26^{\prime}28^{\prime\prime}\!.4$
-Surface dipole magnetic field, $B_{s}$ $9.8\times 10^{10}$ G
-Spin-down luminosity, $\dot{E}$ $1.1\times 10^{31}$ erg s-1
-Characteristic age, $\tau_{c}\equiv P/2\dot{P}$ 303 Myr
-Pre-glitch Timing Solution (2002–2014)
-Epoch of ephemeris (MJD TDB) 54547.00000198
-Span of ephemeris (MJD) 52266–56829
-Frequency, $f$ 2.357763492491(28) s-1
-Frequency derivative, $\dot{f}$ $1.2317(66)\times 10^{-16}$ s-2
-Period, $P$ 0.4241307506816(50) s
-Period derivative, $\dot{P}$ $2.216(12)\times 10^{-17}$
-$\chi^{2}_{\nu}[{\rm DoF}]$ 1.80[25]
-Post-Glitch Timing Solution (2016–2023)
-Epoch of ephemeris (MJD TDB) 58144.00000219
-Span of ephemeris (MJD) 57597–59986
-Frequency, $f$ 2.35776345898(14) s-1
-Frequency derivative, $\dot{f}$ $1.120(27)\times 10^{-16}$ s-2
-Period, $P$ 0.424130756710(26) s
-Period derivative, $\dot{P}$ $2.015(49)\times 10^{-17}$
-$\chi^{2}_{\nu}[{\rm DoF}]$ 1.26[35]
-Glitch epoch (MJD) 57041bbEpoch of the glitch estimated by matching the frequency of the two timing solutions; this assumes a constant post-glitch $\dot{f}$.
-$\Delta f$ $3.87(39)\times 10^{-9}$ s-1
-$\Delta f/f_{\rm pred}$ $1.64(16)\times 10^{-9}$
-$\Delta\dot{f}$ $1.12(28)\times 10^{-17}$ s-2
-$\Delta\dot{f}/\dot{f}$ 0.091(23)-0.091(23)
-Table 3: Alternative Timing Solutions for 1E 1207.4-5209
-[…]
-Quadratic Timing Solution (2000–2023)
-Epoch of ephemeris (MJD TDB) 55478.00000467
-Span of ephemeris (MJD) 51549–59986
-Frequency, $f$ 2.357763483778(12) s-1
-Frequency derivative, $\dot{f}$ $1.1128(15)\times 10^{-16}$ s-2
-Period, $P$ 0.4241307522490(21) s
-Period derivative, $\dot{P}$ $2.0018(28)\times 10^{-17}$
-$\chi^{2}_{\nu}[{\rm DoF}]$ 8.55[64]
-Cubic Timing Solution (2000–2023)
-Epoch of ephemeris (MJD TDB) 55478.00000453
-Span of ephemeris (MJD) 51549–59986
-Frequency, $f$ 2.357763483219(32) s-1
-Frequency derivative, $\dot{f}$ $1.1260(17)\times 10^{-16}$ s-2
-Frequency second derivative, $\ddot{f}$ $4.67(25)\times 10^{-26}$ s-3
-Period, $P$ 0.4241307523495(58) s
-Period derivative, $\dot{P}$ $2.0255(30)\times 10^{-17}$
-Period second derivative, $\ddot{P}$ $8.40(45)\times 10^{-27}$ s-1
-$\chi^{2}_{\nu}[{\rm DoF}]$ 3.17[63]
-Cubic Timing Solution (2002–2023)
-Epoch of ephemeris (MJD TDB) 55478.00000447
-Span of ephemeris (MJD) 52266–59986
-Frequency, $f$ 2.357763483142(35) s-1
-Frequency derivative, $\dot{f}$ $1.1318(20)\times 10^{-16}$ s-2
-Frequency second derivative, $\ddot{f}$ $5.51(29)\times 10^{-26}$ s-3
-Period, $P$ 0.42413075236(63) s
-Period derivative, $\dot{P}$ $2.0359(36)\times 10^{-17}$
-Period second derivative, $\ddot{P}$ $0.92(53)\times 10^{-26}$ s-1
-$\chi^{2}_{\nu}[{\rm DoF}]$ 2.72[62]
-Figure 1: Glitch model for the spin-down of 1E 1207.4-5209.
-Top: pulse-phase residuals from the pre-glitch timing solution presented in Table 2.
-The glitch epoch of 2015 January 19 is estimated by matching the frequency of the pre- and post-glitch solutions.
-The year 2000 Chandra data point is not included in the fit (see Section 2.1 for details).
-Bottom: combined residuals from fits to independent timing models for the pre- and post-glitch intervals, respectively.
-The overall $\chi^{2}_{\nu}$ statistic for the fit is 1.49 for 60 DoF, taking into account the fit parameters for each interval.
-Figure 2: Pulse-phase residuals for 1E 1207.4-5209 using the alternative timing models presented in Table 3 that do not involve a glitch.
-These fits include the year 2000 Chandra data point (see Section 2.1 for details).
-Top: A quadratic model fit leaves a sinusoidal-like oscillation in the phase residuals.
-Bottom: the residuals from a cubic model fit (including the frequency second derivative).
-Figure 1 graphs the ToA phase residuals from the pre-glitch timing solution given in Table 2, developed using data points from 2002–2014 (as in Gotthelf & Halpern, 2020).
-As noted in our earlier work, the very first ToA, the Chandra observation of 2000, does not appear consistent with the pre-glitch timing solution, and may indicate an even earlier glitch.
-We again excluded this data point from the fit, but continue to plot it in Figure 1 for reference.
-After 2015, deviations from the expected phase of this solution were evident, and we were no longer able to extend the phase-connected solution forward in time.
-Instead, we postulated a glitch, and found that the subsequent ToAs could be characterized by a new, post-glitch phase-connected ephemeris.
-In this work, we have extended the post-glitch ephemeris, which remains consistent with the earlier result and shows no evidence of a newer glitch.
-The updated change of frequency between the post- and pre-glitch timing solutions is $\Delta f=(3.87\pm 0.39)\times 10^{-9}$ Hz, with a glitch magnitude of $\Delta f/f_{\rm pred}=(1.64\pm 0.16)\times 10^{-9}$.
-The post-glitch solution matches $f_{\rm pred}$, the frequency predicted by the pre-glitch solution on 2015 January 19 (MJD 57041), which we estimate as the epoch of the glitch (ignoring any short-timescale recovery behavior, which would not have been sampled).
-Of note, the glitch magnitude for 1E 1207.4-5209 is typical of rotation-powered pulsars with strong magnetic fields and rapid spin-down, including the Crab.
-For reference, $\Delta f/f=3\times 10^{-9}$ is where the lower-amplitude peak in the bimodal distribution of glitch magnitudes is centered (Espinoza et al.
-2011).
-The magnitude of the glitch is 23% smaller than the value quoted in Gotthelf & Halpern (2020).
-Nevertheless, this is the same post-glitch timing solution as the one published previously in the sense that the cycle counts calculated over the common time span are the same.
-The fit statistic of the combined pre- and post-glitch data sets is $\chi^{2}_{\nu}=1.49$ for 60 DoF.
-This is to be compared to $\chi^{2}_{\nu}=1.44$ for 42 DoF reported previously for the shorter span.
-The uncertainty on the frequency derivative of the revised post-glitch timing solution has grown, perhaps an effect of timing noise.
-In addition, the post-glitch frequency derivative is larger (less negative) than the pre-glitch value by $\approx 9\%$, albeit at the $4\sigma$ level.
-This change is much larger than the $\Delta\dot{f}/\dot{f}\sim 10^{-3}$ typically seen across glitches, which also suggests that timing noise may predominate.
-Weak glitches can be difficult to detect in the presence of timing noise, and timing noise can masquerade as glitches (Antonopoulou et al., 2022).
-Following Gotthelf & Halpern (2020), we therefore considered alternative timing models that do not require the use of a glitch to characterize the timing behavior of 1E 1207.4-5209.
-First, we show that a fit to the data using a quadratic model only, parameterized by the frequency and first derivative, and including the year 2000 observation, leaves a strong sinusoidal-like residual in the phases, with an amplitude of $\sim 0.1$ cycles (see Figure 2).
-As in Gotthelf & Halpern (2020), this results in an unacceptable fit statistic of $\chi^{2}_{\nu}=8.549$ for 64 DoF, substantially worse than for the glitch model.
-The sinusoidal residual with a period similar to the data span suggests that a polynomial term of higher degree should be added to the model.
-The lower panel of Figure 2 shows the result of fitting a cubic model that includes a frequency second derivative.
-This reduces the statistic to $\chi^{2}_{\nu}=3.17$ for 63 DoF, still significantly higher than for the glitch model.
-By excluding the 2000 data point, the fit statistic improves marginally to $\chi^{2}_{\nu}=2.72$ for 62 DoF, but still suggests larger than expected timing noise, as will be discussed in Section 3.
-Parameters from the quadratic and cubic fits are given in Table 3.
-It is important that each of these models, as well as the glitch model, have the same cycle count between all adjacent ToAs.
-Thus they are all the same phase-connected solution that we believe is the true one, not an alias.
-Finally, we note that the binary orbit model suggested in Gotthelf & Halpern (2020) remains a viable interpretation, although less likely a priori.
-2.2 PSR J0821-4300
-Table 4 is the log of timing observations of PSR J0821-4300 used in this work.
-These include the Chandra and XMM-Newton data sets previously reported in Gotthelf et al.
-(2013) that span years 200920122009-2012, but exclude the two observations from 2001.
-In retrospect, it is not possible to span the large time gap from 2001 to 2009 due to the uncertainty in extrapolation of the pre-glitch ephemeris.
-In creating ToAs for PSR J0821-4300, we restrict the extracted photon energies to the 1.54.51.5-4.5 keV range because of the unique energy dependence of the phase of the pulse.
-Below 1.5 keV, the pulse is shifted $\approx 180^{\circ}$, which effectively cancels the pulsations if the full energy band is used (see Gotthelf & Halpern, 2009, for details).
-Photon arrival times were extracted using $30^{\prime\prime}$ and $1\farcs 2$ radius circular apertures for the XMM-Newton and Chandra data sets, respectively.
-Following the results presented in Gotthelf et al.
-(2013), we were able to continue monitoring this pulsar and extend its ephemeris, but only up to 2014.
-Subsequently, the yearly monitoring showed increasing departures of the pulse arrival times from the refined ephemeris.
-Motivated by the similarity to the glitch-like behavior of 1E 1207.4-5209, we then initiated a renewed joint XMM-Newton and Chandra timing campaign on PSR J0821-4300 in 201920202019-2020 to generate an independent phase-connected timing solution for comparison.
-In Figure 3, we postulate that a glitch in the pulse arrival times most likely occurred between the 2015 and 2016 observations.
-Indeed, the monotonic deviations from the pre-glitch timing solution given in Table 5 imply a change in frequency of magnitude $\Delta f/f_{\rm pred}=7.1(6)\times 10^{-10}$, a value similar to that found for 1E 1207.4-5209.
-Considering the pre- and post-glitch timing solutions, the statistic for the combined data set is $\chi^{2}_{\nu}=2.20$ for 32 DoF, taking into account the fit parameters for each interval.
-The sparse coverage around the time of the glitch, between 2014 and 2017, prevents a definitive description of the glitch behavior.
-However, the formal fits show curvature in the trend of the post-glitch residuals relative to the pre-glitch solution, corresponding to a $\approx 7\%$ change in frequency derivative, becoming less negative, similar to the case of 1E 1207.4-5209 but at a higher $9\sigma$ level of significance (see Table 5).
-Curiously, a model involving an anti-glitch (decrease in frequency) fits nearly as a well as the glitch model, with the event occurring at an earlier time, around 2011 September (Figure 4).
-This model parameters are given in Table 6, where it can be seen that magnitude of the anti-glitch is similar to that in the glitch model, while the statistic for the combined data set is slightly worse at $\chi^{2}_{\nu}=2.77$.
-Since the pre-glitch segment in this model is very short, the initial frequency derivative was only measured to a precision of $\approx 8\%$, but it is consistent with the post-glitch value.
-Note that this model has the same set of cycle counts as the glitch model, despite being a different analytic and qualitative description of the same timing data.
-We also consider alternative timing models without the use of a glitch to characterize the timing of PSR J0821-4300.
-A basic quadratic model leaves phase residuals with $\sim 0.1$ cycle amplitude, similar in magnitude to those of 1E 1207.4-5209, but not clearly sinusoidal in nature (Figure 5).
-The fit statistic is $\chi^{2}_{\nu}=6.79$ for 35 DoF, again substantially worse than for the glitch model and similar to that found for 1E 1207.4-5209.
-The additional of a second frequency derivative term for a cubic model has little effect on the shape of the residual curve, and only reduces $\chi^{2}_{\nu}$ to 5.80 for 34 DoF.
-Parameters from the quadratic and cubic fits are given in Table 7.
-It can be seen how the residual curve in Figure 5 could also be interpreted as an early anti-glitch or a later glitch.
-Again, none of these are distinct timing solutions (aliases); they all have the same set of cycle counts.
 Table 8: Log of X-ray Timing Observations of PSR J0821-4300
-Mission Instrument ObsID Date ExpaaExposure times for the XMM-Newton EPIC-pn do not reflect the 29% deadtime in SmallWindow (sw) mode.
-/Mode (UT) (ks)
+Mission Instrument/Mode ObsID Date ExpaaExposure times for the XMM-Newton EPIC-pn do not reflect the 29% deadtime in SmallWindow (sw) mode.
+(UT) (ks)
 XMM EPIC-pn/sw 0606280101 2009 Dec 17, 18 85
@@
 Chandra ACIS-S3/$C$ 22676, 26477 2022 Jul 27, 28 52
-Table 5: Glitch Ephemerides for PSR J0821-4300footnotetext: Notes.
-Position and proper motion from Mayer et al.
-(2020); Derived parameters ($B_{s}$, $\dot{E}$, $\tau_{c}$) are based on the pre-glitch timing solution.
-[…]
-Epoch of position and $\mu$ (MJD) 53964.0
-R.A. (J2000) $08^{\rm h}21^{\rm m}57.\!^{\rm s}3653(31)$
-Decl. (J2000) $43^{\circ}00^{\prime}17.\!^{\prime\prime}074(43)$
-R.A. proper motion, $\mu_{\alpha}\,{\rm cos}\,\delta$ $54.1\pm 8.3$ mas yr-1
-Decl. proper motion, $\mu_{\delta}$ $28.1\pm 10.5$ mas yr-1
-Surface dipole magnetic field, $B_{s}$ $2.9\times 10^{10}$ G
-Spin-down luminosity, $\dot{E}$ $1.9\times 10^{32}$ erg s-1
-Characteristic age, $\tau_{c}$ 254 Myr
-Pre-glitch Timing Solution (2010–2014)
-Epoch of ephemeris (MJD TDB) 56065.00000030
-Span of ephemeris (MJD) 55182–56948
-Frequency, $f$ 8.865291025552(85) s-1
-Frequency derivative, $\dot{f}$ $6.801(44)\times 10^{-16}$ s-2
-Period, $P$ 0.1127994554401(11) s
-Period derivative, $\dot{P}$ $8.654(56)\times 10^{-18}$
-$\chi^{2}_{\nu}[{\rm DoF}]$ 2.70[19]
-Post-Glitch Timing Solution (2015–2022)
-Epoch of ephemeris (MJD TDB) 58530.00000022
-Span of ephemeris (MJD) 57272–59787
-Frequency, $f$ 8.865290891814(75) s-1
-Frequency derivative, $\dot{f}$ $6.321(27)\times 10^{-16}$ s-2
-Period, $P$ 0.11279945714171(95) s
-Period derivative, $\dot{P}$ $8.042(34)\times 10^{-18}$
-$\chi^{2}_{\nu}[{\rm DoF}]$ 1.47[13]
-Glitch epoch (MJD) 57362bbEpoch of the glitch estimated by matching the frequency of the two timing solutions; this assumes a constant post-glitch $\dot{f}$.
-$\Delta f$ $6.26(55)\times 10^{-9}$ s-1
-$\Delta f/f_{\rm pred}$ $7.06(62)\times 10^{-10}$
-$\Delta\dot{f}$ $4.80(52)\times 10^{-17}$ s-2
-$\Delta\dot{f}/\dot{f}$ 0.0706(76)-0.0706(76)
-Table 6: Anti-glitch Ephemerides for PSR J0821-4300bbfootnotetext: Epoch of the glitch estimated by matching the frequency of the two timing solutions; this assumes a constant post-glitch $\dot{f}$.
-[…]
-Pre-glitch Timing Solution (2010–2011)
-Epoch of ephemeris (MJD TDB) 55484.00000014
-Span of ephemeris (MJD) 55182–55784
-Frequency, $f$ 8.86529106058(42) s-1
-Frequency derivative, $\dot{f}$ $6.14(50)\times 10^{-16}$ s-2
-Period, $P$ 0.1127994549943(53) s
-Period derivative, $\dot{P}$ $7.81(64)\times 10^{-18}$
-$\chi^{2}_{\nu}[{\rm DoF}]$ 2.535[11]
-Post-Glitch Timing Solution (2012–2022)
-Epoch of ephemeris (MJD TDB) 57830.00000060
-Span of ephemeris (MJD) 55873–59787
-Frequency, $f$ 8.865290929387(33) s-1
-Frequency derivative, $\dot{f}$ $6.2179(81)\times 10^{-16}$ s-2
-Period, $P$ 0.11279945666364(42) s
-Period derivative, $\dot{P}$ $7.912(10)\times 10^{-18}$
-$\chi^{2}_{\nu}[{\rm DoF}]$ 2.695[21]
-Glitch epoch (MJD) 55828b
-$\Delta f$ $5.4(6)\times 10^{-9}$ s-1
-$\Delta f/f_{\rm pred}$ $6.1(6)\times 10^{-10}$
-Table 7: Alternative Timing Solutions for PSR J0821-4300
-[…]
-Quadratic Timing Solution (2010–2022)
-Epoch of ephemeris (MJD TDB) 57485.00000096
-Span of ephemeris (MJD) 55182–59787
-Frequency, $f$ 8.865290948152(24) s-1
-Frequency derivative, $\dot{f}$ $6.2723(51)\times 10^{-16}$ s-2
-Period, $P$ 0.11279945642488(30) s
-Period derivative, $\dot{P}$ $7.9807(65)\times 10^{-18}$
-$\chi^{2}_{\nu}[{\rm DoF}]$ 6.79[35]
-Cubic Timing Solution (2010–2022)
-Epoch of ephemeris (MJD TDB) 57485.00000097
-Span of ephemeris (MJD) 55182–59787
-Frequency, $f$ 8.865290947727(71) s-1
-Frequency derivative, $\dot{f}$ $6.2665(52)\times 10^{-16}$ s-2
-Frequency second derivative, $\ddot{f}$ $1.00(16)\times 10^{-25}$ s-3
-Period, $P$ 0.11279945643030(91) s
-Period derivative, $\dot{P}$ $7.9733(66)\times 10^{-18}$
-Period second derivative, $\ddot{P}$ $1.27(20)\times 10^{-27}$ s-1
-$\chi^{2}_{\nu}[{\rm DoF}]$ 5.81[34]
-Figure 3: Glitch model for the spin-down of PSR J0821-4300.
-Top: pulse-phase residuals from the pre-glitch timing solution presented in Table 5.
-The glitch epoch of 2015 December 6 is estimated by matching the frequency of the pre- and post-glitch solutions.
-Bottom: combined residuals from fits to independent timing models for the pre- and post-glitch intervals, respectively.
-The overall $\chi^{2}_{\nu}$ statistic for the fit is 2.20 for 32 DoF, taking into account the fit parameters for each interval.
-Figure 4: Anti-glitch model for the spin-down of PSR J0821-4300.
-Top: pulse-phase residuals from the pre-glitch timing solution presented in Table 6.
-The glitch epoch of 2012 March 13 is estimated by matching the frequency of the pre- and post-glitch solutions.
-Bottom: combined residuals from fits to independent timing models for the pre- and post-glitch intervals, respectively.
-The overall $\chi^{2}_{\nu}$ statistic for the fit is 2.77 for 32 DoF, taking into account the fit parameters for each interval.
-Figure 5: Pulse-phase residuals for PSR J0821-4300 using the alternative timing models presented in Table 7 that do not involve a glitch.
-Top: a quadratic model.
-Bottom: a cubic model (including the frequency second derivative).
-2.3 PSR J1852++0040
-PSR J1852++0040 was the first CCO to have its spin-down rate measured (Halpern & Gotthelf, 2010), but there followed a 12 yr hiatus in its monitoring.
-We have obtained a new set of observations of PSR J1852++0040 from a joint Chandra and XMM-Newton investigation designed to restart a phase-connected timing solution in 2021–2023 (Table 8).
-Photon arrival times were extracted in the 151-5 keV energy range using a $12^{\prime\prime}$ or $18^{\prime\prime}$ radius aperture with XMM-Newton, and a $1\farcs 8$ radius aperture with Chandra.
-Table 9 presents the new results from the 10 observations obtained in 2021–2023, along with results from the 2004–2009 era (Halpern & Gotthelf, 2010).
-As shown in Figure 6, we are able to obtain independent timing solutions for PSR J1852++0040 both before and after the 12 yr gap.
-However, the long gap does not allow the cycle count to be maintained, given the uncertainty on the frequency derivative of the earlier solution.
-During the 2021–2023 epoch, the frequency derivative is poorly constrained in the fit, as it contributes less than 2 cycles, based on the prior measured value in 2004–2009.
-The ephemeris for the earlier era predicts the frequency in 2021–2023 to within $\Delta f=4.05\times 10^{-9}$ Hz, comparable to the extrapolated uncertainty on this parameter ($\sigma_{f}=4.03\times\ 10^{-9}$ Hz).
-This is of a similar magnitude as that measured for both the 1E 1207.4-5209 and PSR J0821-4300 glitches.
-Thus, we would not be sensitive to a glitch in PSR J1852++0040 having similar magnitude as inferred for the two other CCO pulsars.
 Table 9: Log of X-ray Timing Observations of PSR J1852++0040
-Mission Instrument ObsID Date ExpaaExposure times for the XMM-Newton EPIC-pn do not reflect the 29% deadtime in SmallWindow (sw) mode.
-/Mode (UT) (ks)
+Mission Instrument/Mode ObsID Date ExpaaExposure times for the XMM-Newton EPIC-pn do not reflect the 29% deadtime in SmallWindow (sw) mode.
+(UT) (ks)
 XMM EPIC-pn/sw 0204970201 2004 Oct 18 30.6
@@
 Chandra ACIS-S3/$C$ 27823 2023 May 2 31.8
-Table 9: Timing Solutions for PSR J1852++0040footnotetext: Notes. Derived parameters ($B_{s}$, $\dot{E}$, $\tau_{c}$) are based on the 2004–2009 timing solution.
-[…]
-R.A. (J2000) $18^{\rm h}52^{\rm m}38^{\rm s}\!.57$
-Decl. (J2000) $00^{\circ}40^{\prime}19^{\prime\prime}\!.8$
-Surface dipole magnetic field, $B_{s}$ $3.1\times 10^{10}$ G
-Spin-down luminosity, $\dot{E}$ $3.0\times 10^{32}$ erg s-1
-Characteristic age, $\tau_{c}$ 192 Myr
-Quadratic Timing Solution (2004–2009)
-Epoch of ephemeris (MJD TDB) 54168.00000093
-Span of ephemeris (MJD) 53296–55041
-Frequency, $f$ 9.53174258208(17) s-1
-Frequency derivative, $\dot{f}$ $7.882(81)\times 10^{-16}$ s-2
-Period, $P$ 0.1049126108253(19) s
-Period derivative, $\dot{P}$ $8.676(90)\times 10^{-18}$
-$\chi^{2}_{\nu}[{\rm DoF}]$ 1.01[20]
-Quadratic Timing Solution (2021–2023)
-Epoch of ephemeris (MJD TDB) 59677.00000016
-Span of ephemeris (MJD) 59288–60066
-Frequency, $f$ 9.53174220286(34) s-1
-Frequency derivative, $\dot{f}$ $8.62(45)\times 10^{-16}$ s-2
-Period, $P$ 0.1049126149991(37) s
-Period derivative, $\dot{P}$ $9.48(50)\times 10^{-18}$
-$\chi^{2}_{\nu}[{\rm DoF}]$ 0.976[7]
-3 Discussion
-3.1 Do CCOs Glitch?
-New timing observations reported here extend our previous ephemeris for 1E 1207.4-5209, (Gotthelf & Halpern, 2020) and reveal similar noisy timing behavior in PSR J0821-4300.
-However, because of the sparse sampling, and the relatively imprecise ToAs of thermal X-ray pulses compared with those of radio pulsars, these results do not conclusively prove that glitches occur in CCOs.
-A glitch is a priori surprising in a CCO, as discussed in (Gotthelf & Halpern, 2020), because CCOs have $|\dot{f}|<10^{-15}$ s-2.
-The glitch activity measured in pulsars is best correlated with $\dot{f}$, while it is rare that any pulsar with $|\dot{f}|<10^{-15}$ s-2 is seen to glitch.
-According to Antonopoulou et al.
-(2022), PSR B0410+69 is the only pulsar with $\dot{f}$ as small as $\approx-5\times 10^{-16}$ s-2 that was observed to glitch.
-Ultimately, there is a limit in which a weak glitch can be either obscured or mimicked by timing noise, and that limit may have been reached in our study.
-Another manifestation of this ambiguity is the ability of an early anti-glitch to fit the timing of PSR J0821-4300 in our data almost as well as a later glitch does.
-We consider the anti-glitch fit for PSR J0821-4300 to be unconvincing as a physical model because it rests on a short pre-glitch span of data.
-While anti-glitches are sometimes seen in magnetars and accreting pulsars, they are rare in rotation-powered pulsars (Tuo et al., 2024; Panin & Sokolova, 2025).
-In Section 3.2 we discuss timing noise as an alternative to fits involving glitches.
-With these caveats in mind, the implications of the changes in frequency derivative $\dot{f}$ required in our glitch fits deserve explicit mention.
-The newly extended post-glitch timing solution for 1E 1207.4-5209, by refining its frequency derivative from the previous study, revealed a large relative change of $\approx 9\%$ in $\dot{f}$ across the putative glitch.
-That is, $\Delta\dot{f}/\dot{f}\approx-0.09$, with $\dot{f}$ post-glitch becoming less negative.
-Note that the magnitude of $\Delta\dot{f}\sim 10^{-17}$ s-2 is itself not unprecedented across glitches (e.g., Figure 11 of Antonopoulou et al.
-2022), but its relative change and sign are unusual.
-These are different from the more common glitches in which $\Delta\dot{f}/\dot{f}\approx 10^{-3}$ with $\dot{f}$ becoming more negative in 80% of the cases (Antonopoulou et al., 2022).
-While the change in $\dot{f}$ in 1E 1207.4-5209 is only a $4\sigma$ result, more definitive evidence of the same effect is found in the case of PSR J0821-4300; its $\dot{f}$ became $7\%$ less negative after the hypothesized glitch, the change being of $9\sigma$ significance.
-If the surface dipole magnetic field strength $B_{s}$ is naively estimated as $\propto f^{-3/2}\dot{f}^{1/2}$, the fitted changes in $\dot{f}$ imply a decrease of $B_{s}$ by a few percent within a decade, a much shorter timescale than expected from the ages of the host SNRs.
-The sign of the change is also in conflict with a leading theory for the absence of old CCOs (Ho, 2011), which postulates that their weak external dipole field $B_{s}$ was caused by prompt burial of a normal $B$ field by supernova fallback debris, while reemergence (increase of $B_{s}$) by ohmic diffusion should follow on timescales of thousands of years (Muslimov & Page, 1995), allowing the CCOs to join the population of canonical pulsars in the $P-\dot{P}$ diagram.
-It should be noted that comparable or larger variations in $\Delta\dot{f}/\dot{f}$ have been discovered from several pulsars in recent years, but they are all associated with changes in magnetospheric emission.
-These phenomena include 1) intermittent pulsars in which the radio pulsations turn off for periods of days to years (Kramer et al., 2006; Camilo et al., 2012; Lyne et al., 2017) while $|\dot{f}|$ decreases by $\geq 50\%$, 2) quasi-periodic changes in radio pulse profile that are associated with a few percent changes in $\dot{f}$ (Lyne et al., 2010), and 3) the radio quiet PSR J2021+4026 that has correlated states of $\gamma$-ray flux and $\dot{f}$, differing by $\approx 20\%$ and $\approx 10\%$, respectively (Allafort et al., 2013; Fiori et al., 2024).
-However, none of these effects are easily applied to CCOs, with their steady, surface thermal X-ray emission and pulse profiles, small spin-down power, and no apparent magnetospheric generation of particles and winds.
-More likely, therefore, the years-long variation of $\dot{f}$ in CCOs has its origin in internal properties that are in common with other young but more energetic pulsars, such as high temperature, strong intrinsic $B$ field and superfluid behavior; in some combination these properties may cause glitches (Ho, 2015) or timing noise.
-3.2 Do CCOs have Unusual Timing Noise?
-Figure 6: Pulse-phase residuals for PSR J1852++0040 from the quadratic spin-down model presented in Table 9, from two independent timing solutions corresponding to the intervals 2004–2009 (top) and 2021–2023 (bottom).
-Fits to the spin-down of 1E 1207.4-5209 and PSR J0821-4300 using only $f$ and $\dot{f}$ without invoking glitches leave systematic trends in their phase residuals that may be regarded as timing noise.
-Here we attempt to evaluate how this compares to timing noise in canonical pulsars with similar spin-down properties.
-Several metrics of timing noise were reviewed by Namkham et al.
-(2019).
-The one most useful here is $\sigma_{\rm TN,2}$ from Shannon & Cordes (2010), in which
-$\sigma_{\rm TN,2}^{2}=\sigma_{\rm R,2}^{2}-\sigma_{\rm W}^{2}.$ (1)
-Here $\sigma_{\rm R,2}$ is the root-mean-square (rms) of the measured residuals from a quadratic fit, and $\sigma_{\rm W}$ is the typical uncertainty of a ToA.
-The subscripts R and W indicate a red or white noise process.
-An advantage of this method is that it explicitly acknowledges the red noise character of timing noise by including in the analysis the total time span $T$ of the observations.
-Shannon & Cordes (2010) found for hundreds of canonical pulsars that the mean value of $\sigma_{\rm TN,2}$ scales with the spin parameters as
-$\bar{\sigma}_{\rm TN,2}=C_{2}\,f^{\alpha}\,|\dot{f}|^{\beta}\,T^{\gamma}\ \ \mu{\rm s},$ (2)
-where $C_{2}=41.7,\alpha=-0.9,\beta=1.00$, and $\gamma=1.9$.
-Recognizing that there is large scatter in $\sigma_{\rm TN,2}$ around a given mean $\bar{\sigma}_{\rm TN,2}$, Shannon & Cordes (2010) modeled its distribution as log-normal, and found that the standard deviation of ln($\sigma_{\rm TN,2}$) = 1.6.
-[…]
-[…]
-Another characterization of timing noise simply uses the frequency second derivative (equivalent to the braking index) added to the fit as an approximate measure of the residuals.
-An example is the cubic fit of 1E 1207.4-5209 in Figure 2, which yields braking index $n=f\ddot{f}\dot{f}^{-2}=8.6\times 10^{6}$.
-In comparison, pulsars with positive $\ddot{f}$ generally have $n<2\times 10^{5}$ (Figure 5 of Namkham et al.
-2019).
-In the case of PSR J0821-4300, it is interesting that fitting a second derivative of $\ddot{f}=1.0\times 10^{-25}$ s-3 barely improves the fit in Figure 5, while the equivalent braking index is $n=2.3\times 10^{5}$.
-For either pulsar, this method indicates large timing noise, but the meaning is not precise as it does not take into account the time span of the observations or the nature of the residuals remaining after the cubic fit.
-For completeness, we recall that an earlier parameterization of timing noise by Arzoumanian et al.
-(1994) used the frequency second derivative measured over a time span of $T=10^{8}$ s to define
-$\Delta_{8}={\rm log}\left({1\over 6f}\,|\ddot{f}|\,T^{3}\right).$
-However, we do not have enough precise ToAs within any $10^{8}$ s time span to measure a significant $\ddot{f}$.
-Finally, we note that the fitted second derivatives representing timing noise are large enough, given the long time spans of the data, to slowly produce the large fractional changes in $\dot{f}$ that, in context of the glitch model, were constrained to be discrete events.
-Neither interpretation of the changes in $\dot{f}$ appears to be clearly preferable, although timing noise is perhaps simpler.
-4 Conclusions
-Continued timing of the known CCO pulsars establishes PSR J0821-4300 as the second one with timing noise or glitches, quantitatively similar to the behavior of 1E 1207.4-5209 established previously.
-We have taken care to ensure that all of the models applied to the timing residuals of a particular pulsar have the same set of cycle counts between observations.
-Thus, they are the same timing solution fitted with different analytic formulas, and not aliases.
-Even so, is difficult to distinguish glitches from timing noise in these CCOs because of the infrequent sampling, and the lower precision of the ToAs that can be obtained from these thermal X-ray pulsations compared to those of typical radio pulsars.
-Nevertheless, whichever description is adopted, the magnitude of the timing irregularities is greater than in rotation-powered pulsars with comparably small spin-down rate and as weak a dipole magnetic field strength as CCOs.
-One quantitative outcome of the extended monitoring is that the spin-down rate $\dot{f}$ in 1E 1207.4-5209 and PSR J0821-4300 appears to slow by $7-9\%$ on a timescale of a decade.
-This is probably not a monotonic trend representing a decrease of the dipole $B$ field, but more likely a representative short-term fluctuation.
-In the absence of evidence for magnetospheric activity and intermittent mode changes such as are seen in certain variable rotation-powered pulsars, the origin of the timing variations in CCOS should be sought in the internal structure of these young NSs, which are likely to have $B$ fields much stronger than their external dipole component.
-Accretion torque noise from a residual fallback disk is even a possible contributor.
-Noisy timing behavior is an important feature of CCOs as a class, and it may inform theories of glitches and timing noise in the general pulsar population.
-Support for this work was provided by NASA through XMM grants 80NSSC18K0452, 80NSSC19K0866, 80NSSC20K1310, 80NSSC21K0819, 80NSSC22K0867, NICER grants 80NSSC19K1461, 80NSSC21K0106, 80NSSC21K1884, 80NSSC22K1300, and Chandra Awards SAO GO7-18063X, SAO GO0-21059X, SAO GO1-22065X issued by the Chandra X-ray Observatory Center, which is operated by the Smithsonian Astrophysical Observatory for and on behalf of NASA under contract NAS8-03060.
-This investigation is based partly on observations obtained with XMM-Newton, an ESA science mission with instruments and contributions directly funded by ESA Member States and NASA.
-This research has made use of data and/or software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC.